### v5

# Timestamp Manipulation: Timestamp-based Nakamoto-style Blockchains are Vulnerable

Abstract
Nakamoto consensuses are the most widely adopted decentralized consensus mechanism in cryptocurrency systems. Since it was proposed in 2008, many studies have focused on analyzing its security. Most of them focus on maximizing the profit of the adversary. Examples include the selfish mining attack and the recent riskless uncle maker (RUM) attack. In this work, we introduce the Staircase-Unrestricted Uncle Maker (SUUM), the first block withholding attack targeting the timestamp-based Nakamoto-style blockchain. Through block withholding, timestamp manipulation, and difficulty risk control, SUUM adversaries are capable of launching persistent attacks with zero cost and minimal difficulty risk characteristics, indefinitely exploiting rewards from honest participants. This creates a self-reinforcing cycle that threatens the security of blockchains. We conduct a comprehensive and systematic evaluation of SUUM, including the attack conditions, its impact on blockchains, and the difficulty risks. We conducted an empirical analysis on three major ETH 1.x-based blockchains, namely Ethereum 1.x, Ethereum Classic, and Ethereum PoW. Statistical results suggest that, as of October 2025, four major mining pools had been launching SUUM attacks by manipulating timestamps. Finally, we further discuss four feasible mitigation measures against SUUM.
I Introduction
Timestamp-based Nakamoto-style Blockchains. Nakamoto-style blockchains, with Bitcoin [1] and Ethereum 1.x [2, 3] as prime examples, have catalyzed the thriving cryptocurrency sector [4]. These systems rely on consensus mechanisms like Proof-of-Work (PoW) to ensure security and consistency [5, 6, 7, 8, 9, 10]. Miners invest computational resources to solve cryptographic puzzles and append new blocks to the chain [11]. The protocol’s incentive mechanism is designed to be fair, rewarding participants in proportion to their contributed computational power [12, 13, 14]. Ethereum 1.x and its popular derivatives (Ethereum PoW [15], Ethereum Fair [16], Ethereum Classic [17]) feature a significantly reduced average block time of 13 seconds compared to Bitcoin’s 10 minutes. This higher throughput, however, leads to frequent soft forks [18], creating opportunities for mining attacks that maliciously manipulate block difficulty to gain an unfair advantage [19, 20, 21, 22, 23, 24, 14].
Existing Attack Models. A significant body of research has focused on identifying and analyzing vulnerabilities in blockchain incentive models. A prominent line of work explores attacks such as selfish mining [25] and its variants. Recently, the Riskless Uncle Maker (RUM) attack [26] demonstrated a novel exploitation vector: timestamp manipulation. Unlike selfish mining that relies on block withholding alone, RUM exploits timestamp rules to inflate block difficulty, achieving risk-free profits. RUM adversaries strategically set block timestamps to artificially inflate their block’s difficulty above that of competing honest blocks, thereby ensuring preferential selection during forks. This attack is proven to be risk-free, incurring no additional overhead.
Limitations and Challenges. The RUM attack constitutes a tangible and impactful threat [26]. Despite this, its threat model is fundamentally constrained by a rigid, risk-averse design. This conservative posture is not a minor oversight but a structural limitation. It directly restricts the attack’s scope and long-term impact, thereby failing to reflect the full spectrum of adversarial behavior in decentralized networks.
Compounding this issue are gaps in how prior work models attack synergy. The existing security landscape thus contains two critical, unaddressed vulnerabilities. Both of these vulnerabilities are directly tied to the practical risks and adversarial incentives detailed in this paper’s subsequent sections:
L1: Constrained Adversarial Capability.
The RUM attack has a core requirement of zero additional difficulty risk [26]. This requirement forces it to rely on an extremely narrow set of attack initiation conditions. These conditions rarely align with the actual block generation dynamics of Ethereum 1.x (we will quantify this later in the simulations in Section VII). Specifically, the RUM attack only triggers under one condition. The condition is the timestamp difference between a newly mined honest block ($B_{1}^{p_{h}}$) and its parent mainchain block ($B_{0}^{p_{h}}$) must fall within the [9, 18) second window. Mathematically, this corresponds to $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor=1$. This constraint is not arbitrary. It ensures the RUM attack avoids any upward fluctuation in the blockchain’s long-term difficulty, thus maintaining its zero-risk feature. Yet this rigidity directly limits the RUM attack’s impact. We will demonstrate this limitation in detail in our subsequent empirical analysis: $\bullet$ Low attack frequency. Ethereum 1.x’s average block time is 13 seconds [2], meaning the [9, 18) second timestamp gap (required for RUM to initiate) occurs in only 12% of honest block pairs (Section VII-A’s steady-state probability calculations confirm this). This caps RUM’s steady-state attack probability at just 1.74% (vs. SUUM’s 16.84%), severely limiting its reward extraction potential. $\bullet$ Capped profitability. RUM’s risk-free constraint prevents it from exploiting scenarios where even minimal difficulty risk could double its attack opportunities. As we later show in Theorem 5, this leaves RUM’s maximum reward share at 26.12% (for the adversary’s relative power $\alpha=0.25$), far below the 33.30% achieved by risk-tolerant strategies like SUUM.
This mismatch between RUM’s assumptions and real-world adversarial incentives. Miners always tolerate tiny, controllable risks for higher returns, which means existing threat models underestimate how adversaries could scale timestamp-based attacks by relaxing risk constraints:
L2: Narrow Attack Vector Integration.
All prior analyses of timestamp manipulation, including those on the RUM attack, treat this technique in isolation. They completely overlook its potential synergy with block withholding. Block withholding is a well-documented strategy in selfish mining [25] and stubborn mining [27]. However, it has never been combined with timestamp exploitation. Such isolated analysis leads to a severe underestimation of attack impact. We will later prove this through the SUUM attack (Section V) and validate it via simulations (Section VII-D): $\bullet$ Single vs. cascading harm. The RUM attack uses timestamp manipulation alone. Each attack can only replace one honest block, turning it into an uncle block. However, when combined with block withholding, the situation changes. Adversaries first build a private chain using withheld blocks. They then release these blocks. In this case, timestamp manipulation ensures every block in the private chain has higher difficulty than honest blocks at the same height. As shown in Figure 3-(c), this results in cascading chain reorganizations. A single SUUM attack can invalidate multiple honest blocks compared to only one block invalidated by RUM. When the adversary’s relative power $\alpha=0.3$, honest rewards decrease by 11.85% (Section VII-B). $\bullet$ Loss of persistence. The RUM attack relies on an isolated strategy. It cannot sustain attacks over the long term. The reason is that it lacks a mechanism to lock in advantages. By contrast, the SUUM attack uses block withholding to create a self-reinforcing cycle. Withheld blocks with manipulated timestamps suppress the growth of honest blocks’ difficulty (Section V-C). At the same time, chain reorganizations allow the adversary to steal rewards. This makes it economically rational for honest miners to join the attack. We formalize this death spiral in Theorem 8.
Failure to model such synergy leads to a critical issue. Current threat models cannot capture the systemic and permanent harm. This harm is inflicted on the blockchain’s economic model by coordinated strategies. We address this gap through SUUM’s integrated framework. The above limitations leave fundamental questions unresolved:
Q1: Can RUM adversaries systematically relax the constraints to achieve higher rewards and greater persistence?
Q2: Does there exist a synergistic attack strategy that combines timestamp manipulation with block withholding to inflict permanent, systemic harm on the protocol’s economic model?
Proposed Attack Frameworks. To address these questions, we propose two advanced attack strategies that fundamentally extend the threat model for timestamp-based Nakamoto-style blockchains:
S1 to Q1: Unrestricted Uncle Maker (UUM) Attack.
The UUM attack strategically abandons the zero-cost constraint of RUM, expanding the scope of attack initiation conditions to achieve higher returns. By tolerating minimal, carefully calibrated difficulty risk, UUM adversaries significantly increase their steady-state probability of being in an attack state, thereby amplifying their reward share compared to RUM.
S2 to Q2: Staircase-Unrestricted Uncle Maker (SUUM) Attack.
As a strategic escalation of UUM, the SUUM attack integrates block withholding and release mechanisms with precision timestamp manipulation. This synergy enables persistent, near zero-cost attacks. SUUM forms a self-reinforcing cycle: it suppresses honest participants’ returns by manipulating difficulty growth rates while amplifying adversarial profits through strategic chain reorganizations. We formalize SUUM via a three-pillar framework: 1) timestamp manipulation for difficulty dominance, 2) block withholding for reorganization cascades, and 3) difficulty risk control via temporal gap optimization.
UUM Attack. The UUM attack operationalizes a three-phase framework to achieve low-risk, high-reward objectives by capitalizing on timestamp manipulation within Ethereum 1.x-style fork selection rules. In the Deployment Phase, the adversary monitors the timestamp difference between a newly generated honest block and its preceding mainchain block, waiting for this difference to be greater than or equal to 9 seconds. This relaxed condition, compared to RUM’s narrower window, significantly increases the frequency of attack opportunities. Once the condition is met, the system enters the Execution Phase. Here, the adversary mines on the parent block and, upon finding a block before honest participants, sets its timestamp to create a difficulty advantage. Specifically, the adversary ensures the timestamp difference for its block is between 1 and 900 seconds, making its difficulty higher than the competing honest block and triggering a successful fork.
In the Minimum Risk Control Phase, the adversary meticulously calibrates the timestamp to maximize the difficulty advantage while minimizing the subsequent impact on the blockchain’s long-term difficulty growth, often by setting the floor function of the timestamp difference to 1. This phased approach allows UUM to trade a minimal, controlled amount of risk for a substantially higher steady-state attack probability and reward yield than RUM.
Figure 1: Attack Flowchart. This figure illustrates the attack flowchart for the proposed RUM, UUM and SUUM attacks in timestamp-based Nakamoto-style blockchains. The flowchart delineates the systematic process through which adversaries manipulate block timestamps and strategically withhold or release blocks to gain disproportionate rewards. It highlights the adversarial strategies’ escalation from RUM (risk-free) to UUM (risk-tolerant) and SUUM (withholding-enabled), emphasizing their impact on blockchain protocol and incentive fairness.
SUUM Attack. Building on UUM’s risk-tolerant design, SUUM further integrates block withholding to enable persistent, cost-free attacks. It is orchestrated through a sophisticated five-phase model. The attack begins in the SUUM Deployment Phase. If the adversary finds a block first, it withholds it, transitioning to the SUUM Block Withholding Phase and building a private chain. This mirrors selfish mining [25] but is augmented with precise timestamp control. If an honest miner finds a block that meets certain timestamp conditions (timestamp difference $\geq$ 9 seconds), SUUM can Downgrade to a UUM-like strategy, ensuring it remains profitable even when not leading.
The core of SUUM’s power lies in the SUUM Releasing Block Phase. When an honest block is published, the adversary immediately releases one or more withheld blocks from its private chain. Crucially, each released block has its timestamp manipulated to ensure its difficulty exceeds that of the honest chain at the same height, causing a chain reorganization that turns honest blocks into uncles or orphans. This process is governed by the SUUM Minimum Risk Control Phase, where the adversary sets the timestamps of its withheld blocks to be just high enough to win fork competitions while minimizing the resultant difficulty increase, thus preserving its future mining efficiency. This combination of withholding and timestamp manipulation creates a staircase effect, where the adversary can repeatedly release blocks to reorganize the chain, maximizing rewards and inflicting sustained damage on honest miners’ profitability.
Our Results. Through extensive discrete-event simulations of a timestamp-based Nakamoto-style blockchain, we empirically validate that SUUM achieves superior reward extraction (e.g., 33.30% vs. 28.41% for UUM and 26.12% for RUM at adversary’s relative power $\alpha=0.25$) and induces a higher forking rate (16.84% vs. 13.40% for UUM and 1.74% for RUM). Furthermore, we demonstrate that SUUM triggers a death spiral in the protocol’s economic model, where rational miners are incentivized to join adversarial coalitions, accelerating protocol abandonment.
Latest empirical evidence from mainstream ETH 1.x-style blockchains further underscores the urgency of timestamp manipulation threats. We analyzed 50,130 mainchain blocks across three prominent networks: Ethereum 1.x (block heights 15505647–15535776), Ethereum Classic (23232147–23242146), and Ethereum PoW (22905613–22915612). We identified the presence of persistent, systematic timestamp manipulation.
Key observations from real-world data across all three blockchains include: 1) The proportion of blocks with timestamp differences divisible by 9 is far lower than the expected normal ratio; 2) The proportion of blocks with timestamp differences equal to multiples of 9 minus 1 is significantly higher than the expected normal ratio; 3) Four major mining pools (Ethereum 1.x: 0x829bd8…, Ethereum PoW: 0x9205c2…, Ethereum Classic: 0x406177… and 0x35aa26…) perfectly avoid generating blocks with timestamp differences that are multiples of 9. This deliberate timestamp manipulation aims to increase the probability of winning in fork competitions and avoid block difficulty fluctuations, which highly aligns with the minimal risk control strategy (Theorem 7) in our proposed SUUM/SUM attacks. These findings confirm that timestamp manipulation is not a transient issue limited to the history of Ethereum 1.x, but a persistent and widespread threat across all ETH 1.x-style blockchains.
Figure 2: State Transition Process. This figure illustrates the state transition process under different mining strategies. It delineates the dynamic transformation relationships among different states in the blockchain system. Black + Cyan transitions denote RUM [26]. Black + Violet transitions denote UUM. Red transitions denote SUUM. It clearly presents the path from the initial state to the attack state through nodes and arrows, including the critical transition conditions between the deployment state and the attack state. It annotates the probabilities of state transitions and the behaviors of participants, revealing how attack strategies influence the blockchain’s difficulty and reward distribution by manipulating timestamps and block release timing. (a) In the RUM attack, the adversary employs a risk-free strategy and is not allowed to withhold blocks. The risk-free condition is achieved through the transition from the deployment state to the attack state when $t_{1}^{p_{h}}-t_{0}^{p_{h}}\in[9,18)$. (b) Building upon the RUM framework, the UUM attack amplifies the steady-state probability of transitioning to the attack state, trading minimal risk for higher rewards. Specifically, the condition for transitioning from the deployment state to the attack state is extended to $t_{1}^{(p_{h})}-t_{0}^{(p_{h})}\in[9,+\infty)$. During the attack state, once the UUM adversary successfully discovers a new block, the attack is deemed successful by strategically setting the timestamp $t_{1}^{(p_{\alpha})}-t_{0}^{(p_{\alpha})}\in[1,900)$. (c) Building upon the UUM framework, the SUUM attack further enhances the adversary’s strategy space by allowing the withholding of blocks. The Markov process of SUUM can be decomposed into two parts. On one hand, from the local perspective of the deployment and downgrade states, SUUM reduces to UUM, where the strategies of the two adversaries are identical. On the other hand, from the local perspective of the deployment and attack states, SUUM resembles traditional selfish mining, with the sole distinction being that the SUUM adversary meticulously manipulates the timestamp $t_{i}^{(p_{h})}-t_{i}^{(p_{\alpha})}\in[1,9)$ to ensure the success of the SUUM attack.
Our Contributions. We summarize the contributions of this paper as follows:
  1. $\bullet$
    Synergistic Attack Vector Design. We formalize a three-pillar framework combining 1) timestamp manipulation to inflate difficulty, 2) block withholding to maximize reorgs, and 3) difficulty risk control via temporal gap optimization. This synergy enables adversaries to systematically drain rewards while suppressing honest participation.
  2. $\bullet$
    Permanent Protocol Harm. The proposed UUM/SUUM attacks induce permanent inflation of block difficulty, creating persistent reward distortions that endure even post-attack, thereby eroding the protocol’s security.
  3. $\bullet$
    Cost-free Attack Sustainability. SUUM adversaries achieve cost-free persistence through second-level timestamp manipulation and strategic block withholding, eliminating traditional constraints like hash power thresholds. Simulations confirm minimal difficulty escalation (0.21 maximal risk) despite sustained exploitation.
  4. $\bullet$
    Empirical Validation. We implement a discrete-event timestamp-based Nakamoto-style blockchan simulator and conduct large-scale experiments (1M blocks, 10K trials) to quantify SUUM’s dominance: adversaries with $\alpha=0.3$ reduce honest rewards by 11.85%, while SUUM’s forking rate (16.84%) exceeds UUM (13.40%) and RUM (1.74%) at equivalent power.
  5. $\bullet$
    Exposing Timestamp Manipulation Attack Pools. Through an investigation of three major Ethereum 1.x-style blockchains, we confirm that timestamp manipulation attacks have been implemented in real-world environments, and expose four mining pools that are currently executing such attacks.
  6. $\bullet$
    Mitigation Framework. We propose countermeasures including timestamp consensus mechanisms and difficulty decoupling protocols, establishing defense principles against timestamp manipulation attacks.
II Threat Model
Participant Composition. We assume a decentralized protocol, denoted as $\Gamma$, which is collectively executed by a set of participants, $\mathcal{P}$, across consecutive time slots. The set of participants, $\mathcal{P}$, can be subdivided into two subsets: the set of honest participants, $\mathcal{P}_{\mathcal{H}}=\left\{p_{h}^{1},~p_{h}^{2},\ldots,p_{h}^{n}\right\}$, who strictly adhere to the protocol $\Gamma$, and the set of adversaries, $\mathcal{P}_{\mathcal{A}}$. Consequently, the entire set $\mathcal{P}$ can be represented as the union of $\mathcal{P}_{\mathcal{H}}$ and $\mathcal{P}_{\mathcal{A}}$, i.e., $\mathcal{P}=\mathcal{P}_{\mathcal{H}}\cup\mathcal{P}_{\mathcal{A}}$.
It is noteworthy that the scope of honest mining participants, $\mathcal{P}_{\mathcal{H}}$, is not limited to a single entity. Its composition can be extended to mining pool organizations and even consortia formed by multiple mining pools. Correspondingly, the adversary set, $\mathcal{P}_{\mathcal{A}}$, may also exist in the form of mining pools or mining pool alliances, exhibiting considerable complexity and diversity. This paper follows the common assumptions in the blockchain literature [25, 2, 7], namely, that during the attack period, the hash power of each participant remains constant, no new participants join the network, all mining hardware is preconfigured, and relevant cost (including but not limited to electricity expenses and mining hardware acquisition cost) are prepaid. In practice, historical data indicate that the active hashing rate in the network remains relatively stable over short periods, with minor fluctuations [28].
In protocol $\Gamma$, each participant $p\in\mathcal{P}$ is associated with a numerical value $\mu_{p}$ that lies between 0 and $1$, such that $\sum_{{p}\in\mathcal{P}}\mathbf{\mu}_{{p}}=1$. Here, $\mu_{p}$ represents the participation power ratio of participant $p$ in protocol $\Gamma$, which can be embodied specifically as hash power, stake power, or other relevant metrics.
Regarding the settings of timestamps and block structures, we assume that the timestamp of the genesis block $\mathcal{B}_{0}$ is $t_{0}=0$. Starting from the genesis block, the timestamp of block $\mathcal{B}_{i}$, which is generations of $i$ away from the genesis block on the main chain, is denoted as $t_{i}$, with a corresponding difficulty of $\mathcal{D}_{i}$. Its immediate predecessor (parent block) $\mathcal{B}_{i-1}$ has a timestamp of $t_{i}^{p}=t_{i-1}$. Additionally, a variable ${pu}_{i}\in\left\{0,1\right\}$ is introduced to indicate whether the parent block of block $\mathcal{B}_{i}$ references any uncle blocks. If it does, then ${pu}_{i}=1$; otherwise, ${pu}_{i}=0$.
Adversary’s Target. Ethereum 1.x, as a cryptocurrency system, embraces the same notion of fairness as its counterparts, namely, participants with an expected hash rate proportion of $\mu_{p}$ should be able to mine a corresponding proportion of $\mu_{p}$ blocks and thus obtain $\mu_{p}$ of the mining rewards, as reported in [25, 7]. In this study, the core target of the adversary $\mathcal{P}_{\mathcal{A}}$ is to surpass their fair share based on hash rate, specifically by mining more blocks than their proportional share $\mu_{\mathcal{P}_{\mathcal{A}}}$, thereby obtaining rewards exceeding their deserved portion.
Honest Participant’s Strategy Space. The definition of honest mining protocol in this paper refers to the relevant academic literature in the field of blockchain [25, 29] and specifically follows the rules established by the Ethereum 1.x White Paper [2]. Accordingly, honest participants always strive to mine the blocks with the highest total difficulty and strictly follow the protocol requirements, neither withholding blocks nor engaging in any form of manipulation of block timestamps.
Adversary’s Strategy Space. In contrast, adversaries possess greater freedom in selecting their strategies. Within the framework of this study, adversaries are permitted to deviate from the honest mining protocol to a certain extent, but the blocks they produce must strictly comply with the validity requirements of the Ethereum 1.x protocol rules [2]. Specifically, adversaries are free to adjust the timestamps of the blocks they mine, but must ensure that these timestamps fall within a valid range (as described in Definition 2). Furthermore, adversaries have complete autonomy in choosing which blocks to mine.
To clearly distinguish the attack models proposed in this study from the attack strategies presented in previous literature [30, 31, 25, 29, 26], we introduce two specific types of adversaries: UUM and SUUM. The UUM adversary abandons the risk-free constraint of the RUM attack, increases the steady-state probability of being in the attack state, and possesses the ability to selectively set block timestamps, but is not allowed to withhold blocks. Once a valid block is found by the UUM adversary, it must be immediately released. The SUUM adversary further relaxes the constraints by discarding the rule against withholding blocks in UUM, thereby further increasing its steady-state probability of being in the attack state. Meanwhile, the SUUM adversary retains the right to strategically set block timestamps.
III Review of RUM Attack
In this section, we introduce the RUM attack, which targets vulnerabilities in the incentive mechanism of the Nakamoto-style protocol. The target of this attack is for a RUM adversary who deviates from the protocol to obtain higher rewards than an honest participant.
III-A Preliminaries
Each participant $\mathcal{P}$ in the Proof-of-Work blockchain engages in an iterative process aimed at deriving solutions to the cryptographic puzzle required for constructing a valid block. This process can be rigorously abstracted as a paradigm of Bernoulli trials: participants propose a solution randomly, and if this solution meets the established criteria, the trial outcome is recorded as true; otherwise, it is recorded as false. This series of mutually independent Bernoulli trials collectively constitutes a Bernoulli process. Upon observing this series of trials, the frequency of attempts required to achieve a successful outcome follows a geometric distribution. The duration spent on successfully discovering a valid block is allocated according to an exponential distribution.
Notably, both the geometric and exponential distributions exhibit the property of memorylessness. This implies that the success probability of each trial remains constant and is regulated by the difficulty parameter of the aforementioned protocol. Therefore, the probability of a participant finding a valid solution does not fluctuate due to their previous failures. Once any participant successfully obtains a valid block, by integrating it into their local blockchain and restarting the mining process, their opportunity to mine subsequent blocks remains unchanged. This property is also referred to as progress-free.
The RUM attack exploits the fact of the protocol $\Gamma$ that honest participants $\mathcal{P}_{A}$ default to selecting the block with higher difficulty in the presence of fork competition in Ethereum 1.x. For ease of understanding, we illustrate an example of fork selection through Example 1.
We denote the timestamp and difficulty of the i-th block $\mathcal{B}_{{i}}^{{p}_{{h}}}$ on the honest branch during a fork competition as ${t}_{{i}}^{{p}_{{h}}}$ and $\mathcal{D}_{{i}}^{{p}_{{h}}}$, respectively. Similarly, the timestamp and difficulty of the i-th block $\mathcal{B}_{{i}}^{{p}_{{a}}}$ on the adversary’s branch during a fork competition are denoted as ${t}_{{i}}^{{p}_{{a}}}$ and $\mathcal{D}_{{i}}^{{p}_{{a}}}$.
Example 1 (Fork Selection).
The blockchain comprises three blocks, denoted as $\mathcal{B}_{0}$, $\mathcal{B}_{1}^{{p}_{{a}}}$, and $\mathcal{B}_{1}^{{p}_{{h}}}$. The timestamp of block $\mathcal{B}_{0}$ is represented by ${t}_{0}^{{p}_{{h}}}$ or ${t}_{0}^{{p}_{{a}}}$, and its difficulty is denoted by $\mathcal{D}_{0}^{{p}_{{h}}}$ or $\mathcal{D}_{0}^{{p}_{{a}}}$. For block $\mathcal{B}_{1}^{{p}_{{a}}}$, the timestamp is ${t}_{1}^{{p}_{{a}}}$ and its difficulty is $\mathcal{D}_{1}^{{p}_{{a}}}$. Similarly, for block $\mathcal{B}_{1}^{{p}_{{h}}}$, the timestamp is ${t}_{1}^{{p}_{{h}}}$ and its difficulty is $\mathcal{D}_{1}^{{p}_{{h}}}$. According to the Ethereum 1.x protocol, participants are instructed to select the block with the maximum difficulty. Specifically:
  1. (1)
    Case 1: If $\mathcal{D}_{1}^{{p}_{{h}}}={m}{a}{x}\left\{\mathcal{D}_{1}^{{p}_{{h}}},\mathcal{D}_{1}^{{p}_{{a}}}\right\}$, select block $\mathcal{B}_{1}^{{p}_{{h}}}$.
  2. (2)
    Case 2: If $\mathcal{D}_{1}^{{p}_{{a}}}={m}{a}{x}\left\{\mathcal{D}_{1}^{{p}_{{h}}},\mathcal{D}_{1}^{{p}_{{a}}}\right\}$, select block $\mathcal{B}_{1}^{{p}_{{a}}}$.
  3. (3)
    Case 3: If $\mathcal{D}_{1}^{{p}_{{h}}}=\mathcal{D}_{1}^{{p}_{{a}}}$, select either block arbitrarily.
Now, we provide a formal definition of fairness and block difficulty.
Definition 1 (Fairness).
Let the time be partitioned into discrete intervals $t=1,2,\dots,T$, where each interval $t$ generates $L_{t}$ blocks. A blockchain system satisfies fairness if, for any participant $p_{i}\in\mathcal{P}$, the following condition holds:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\frac{R_{i}^{(t)}}{L_{t}\cdot R}=\frac{1}{T}\sum_{t=1}^{T}\mu_{i}^{(t)},$
where $\mu_{i}^{(t)}$ denotes the power proportion of participant $p_{i}$ during interval $t$.
Definition 1 formally defines the fairness of blockchains. Over the long term, in accordance with the Law of Large Numbers, the reward distribution among participants strictly converges to their resource input (power/stake). This definition not only accounts for the characteristics of blockchain, such as intermittent block generation and resource fluctuations, but also eliminates the interference of short-term random fluctuations via the limit operation. It serves as a core quantitative criterion for evaluating the rationality of incentive mechanisms and whether attacks undermine system fairness.
Definition 2 (Block Difficulty).
The difficulty $\mathcal{D}_{i}$ of block $\mathcal{B}_{i}$ satisfies the following equation:
$\mathcal{D}_{i}\overset{def}{=}{max\left\{{2^{17},\mathcal{D}_{i}^{p}+f\cdot\left\lfloor\frac{\mathcal{D}_{i}^{p}}{2048}\right\rfloor}\right\}},$
where $f={max\left\{{1+{pu}_{i}-\left\lfloor\frac{t_{i}-t_{i-1}}{9}\right\rfloor,-99}\right\}}$.
Please note that in Ethereum 1.x, the difficulty of blocks after height $15$ exceeds $2^{17}$. Consequently, in subsequent analyses concerning block difficulty, the difficulty $\mathcal{D}_{i}$ of block $\mathcal{B}_{i}$ is determined by the following equation:
$\mathcal{D}_{i}\overset{def}{=}\mathcal{D}_{i}^{p}+{max\left\{{1+{pu}_{i}-\left\lfloor\frac{t_{i}-t_{i-1}}{9}\right\rfloor,-99}\right\}}\cdot\left\lfloor\frac{\mathcal{D}_{i}^{p}}{2048}\right\rfloor.$
Each participant $p$ who generates a block has the potential to receive a reward, and different block types correspond to different types of rewards. Below, we formally define block rewards.
Definition 3 (Block Reward).
Blocks are classified into mainchain blocks and non-mainchain blocks.
  1. (1)
    A participant $p$ will receive mining rewards $\mathbb{R}_{c}$, nephew rewards $\mathbb{R}_{n}(\cdot)$, transaction fee rewards $\mathbb{R}_{g}(\cdot)$, and whale rewards $\mathbb{R}_{w}(\cdot)$ when they successfully publish a block $\mathcal{B}_{i}$ with a timestamp $t_{i}$ that is ultimately selected as a mainchain block. The calculation of each type of reward is as follows:
    1. $\bullet$
      $\mathbb{R}_{c}\overset{def}{=}2$;
    2. $\bullet$
      $\mathbb{R}_{n}(m)\overset{def}{=}m\cdot\frac{1}{32}\cdot\mathbb{R}_{c}=m\cdot\frac{1}{16}$, where $m\leq 2$ indicates the number of uncle blocks referenced by block $\mathcal{B}_{i}$;
    3. $\bullet$
      $\mathbb{R}_{g}\left(t_{i}\right)\overset{def}{=}\lambda\cdot\mathrm{\Delta}t_{i}$, where $\lambda$ represents the accumulation rate of transaction fee rewards and $\mathrm{\Delta}t_{i}\overset{def}{=}t_{i}-t_{i}^{p}=t_{i}-t_{i-1}$.
  2. (2)
    A participant $p$ will receive uncle rewards $\mathbb{R}_{u}(\cdot)$ when they successfully publish a block that is ultimately not selected as a mainchain block but is referenced as an uncle block by a mainchain block. The calculation of uncle rewards $\mathbb{R}_{u}(\cdot)$ is as follows:
    1. $\bullet$
      $\mathbb{R}_{u}\left(d_{i}\right)\overset{def}{=}\left\{\begin{matrix}{\frac{8-d_{i}}{8}\cdot 2,}&{d_{i}\in N^{+}~and~1\leq d_{i}\leq 6}\\ {0,}&{otherwise}\end{matrix}\right$., where $d_{i}$ represents the generational distance between the uncle block and the mainchain block that references it.
III-B Method
Based on the previously established rules for calculating block difficulty, it is evident that the difficulty of a block depends on the difficulty of its parent block and the difference in timestamps between the two blocks. An adversary can exploit this rule by meticulously and manually manipulating the block timestamps to make their own block’s difficulty higher than that of blocks on other branches, thereby increasing the likelihood of their block being selected preferentially. The RUM attack leverages this vulnerability. We show the state transition process of RUM in Figure 2-(a).
Specifically, the RUM attack manipulates block timestamps in two phases to create a difficulty advantage and achieve risk-free block prioritization:
  1. $\bullet$
    In the deployment phase, the adversary monitors blocks generated by honest participants and waits for the timestamp difference between a new honest block $\mathcal{B}_{1}^{{p}_{{h}}}$ and the previous mainchain block $\mathcal{B}_{0}^{{p}_{{h}}}$ to fall within the interval $t_{1}^{p_{h}}-t_{0}^{p_{h}}\in(9,18]$ seconds (i.e., $\left\lfloor\frac{t_{1}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor=1$). This condition ensures the honest block is valid and that the adversary’s mining difficulty on the parent block $\mathcal{B}_{0}^{{p}_{{h}}}$ matches the honest mining difficulty on $\mathcal{B}_{1}^{{p}_{{h}}}$. Once this timestamp difference is met, the attack progresses to the execution phase.
  2. $\bullet$
    In the execution phase, the adversary mines on the parent block $\mathcal{B}_{0}^{{p}_{{h}}}$ (identical to $\mathcal{B}_{0}^{{p}_{{h}}}$ and aims to generate a valid block $\mathcal{B}_{1}^{{p}_{{h}}}$ before honest participants. The adversary ensures the timestamp difference between $\mathcal{B}_{1}^{{p}_{{h}}}$ and $\mathcal{B}_{0}^{{p}_{{h}}}$ is $1\leq t_{1}^{p_{a}}-t_{0}^{p_{h}}<9$ seconds (i.e., $\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{h}}}{9}\right\rfloor=0$), making the difficulty of $\mathcal{B}_{1}^{{p}_{{a}}}$ higher than that of the honest block $\mathcal{B}_{1}^{{p}_{{h}}}$. According to Ethereum 1.x’s fork selection rule, which prioritizes the chain with higher difficulty, other participants will prefer the adversary’s block, ensuring attack success. Regardless of success, the process returns to the deployment phase to await the next valid timestamp difference. By precisely controlling timestamp differences, the RUM attack creates a sustained difficulty advantage without additional risk, exploiting the protocol’s rules to secure preferential block inclusion and undermine blockchain fairness.
IV The Design of UUM Attack
On top of the RUM attack, we propose UUM attack. As a strategic extension of the traditional RUM attack, the UUM attack aims to surpass the conservative constraints of the latter by relaxing attack conditions, significantly enhancing the steady-state probability of remaining in the attack state, and maximizing adversaries’ gains.
UUM attackers achieve this by accurately calculating the timestamp difference between the parent block and the current block, leveraging the rule that timestamps influence difficulty calculation to make the difficulty value of their own blocks higher than that of honest nodes’ blocks. This directly triggers the protocol’s fork selection rule (prioritizing the chain with higher difficulty), leading to the discardment of honest blocks by the network. To ensure the long-term covertness of the attack, UUM adversaries meticulously control timestamps to avoid causing abnormal fluctuations in the overall difficulty of the blockchain.
IV-A Method
Specifically, the UUM attack operationalizes a three phase framework to achieve low-risk, high-reward objectives by capitalizing on timestamp manipulation within Ethereum 1.x-style fork selection rules:
  1. $\bullet$
    In the deployment phase, the adversary monitors the timestamp difference between a newly generated honest block $\mathcal{B}_{1}^{{p}_{{h}}}$ and its preceding mainchain block $\mathcal{B}_{0}^{{p}_{{h}}}$, waiting for $t^{p_{h}}_{1}-t^{p_{h}}_{0}\in[9,+\infty)$ (i.e., $\left\lfloor\frac{t^{p_{h}}_{1}-t^{p_{h}}_{0}}{9}\right\rfloor\geq 1$), a condition that validates the honest block and enables selective timestamp manipulation to elevate the difficulty of the adversary’s subsequent block above that of honest competitors. When the floor function result equals 1, the attack reduces to the risk-free RUM variant, whereas values greater than 1 introduce moderate risk alongside enhanced excess rewards.
  2. $\bullet$
    In the execution phase, after satisfying deployment conditions, the adversary mines on the parent block of the latest mainchain block and, upon generating a valid block $\mathcal{B}_{1}^{{p}_{{a}}}$ before honest participants with a timestamp difference $t^{p_{1}}_{a}-t^{p_{0}}_{a}\in[1,900)$, triggers successful attack progression to the risk control phase; this timestamp range ensures block validity and a difficulty advantage ($\mathcal{D}_{1}^{p_{a}}>\mathcal{D}_{0}^{p_{h}}$), thereby compelling honest nodes to prioritize the adversarial block during forks; failed conditions revert the process to deployment.
  3. $\bullet$
    In the minimum risk control phase, the adversary meticulously adjusts the timestamp $t^{p_{1}}_{a}$ of their block to maximize difficulty superiority over honest blocks while minimizing subsequent difficulty growth rates through precise temporal calibration (e.g., enforcing $\left\lfloor\frac{t^{p_{h}}_{1}-t^{p_{a}}_{0}}{9}\right\rfloor=1$ for minimal risk), after which the attack cycle resets to the deployment phase, establishing a self-reinforcing loop of strategic exploitation under controlled difficulty fluctuations.
Next, we continue to elaborate on the relevant initiation condition of the UUM attack under these more relaxed deployment conditions through Theorem 1, further analyzing the advantages and impacts of the UUM attack compared to the RUM attack.
Theorem 1 (UUM Initiation Condition).
The initiation condition for the UUM attack is given by $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$.
Proof of Theorem 1.
The detailed proof of Theorem 1 can be found in Appendix B. ∎
The success of the UUM adversary’s execution phase hinges on the combined effects of multiple critical factors, which not only encompass the difference in block timestamps but are also intimately related to the adversary’s ability to successfully mine blocks that meet specific conditions. To gain an accurate understanding of the characteristics of the UUM attack during its execution phase, we precisely define the conditions for its success through Theorem 2.
Theorem 2 (UUM Successful Condition).
The UUM attack is successful if and only if $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$ and ${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}\in[1,900)$.
Proof of Theorem 2.
The detailed proof of Theorem 2 can be found in Appendix C. ∎
In the course of our in-depth investigation of the UUM attack, we have already explored its initiation condition and success condition. However, for adversaries, minimizing the risks associated with the attack while pursuing success is also a crucial consideration. Here, risk primarily manifests in its impact on the difficulty of the blockchain.
Next, we will elaborate on the specific conditions for minimal risk control in the UUM attack through Theorem 3, further revealing the internal mechanisms and characteristics of the UUM attack under risk control strategies.
Theorem 3 (UUM Successful Condition with Minimal Risk).
The UUM attack is successful with minimal risk if and only if $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor=1$ and ${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}\in[1,900)$.
Proof of Theorem 3.
The detailed proof of Theorem 3 can be found in Appendix D. ∎
As mentioned earlier, the RUM attack has stringent risk control requirements to achieve what is termed a risk-free attack. In contrast, the UUM attack relaxes these constraints to a certain extent, providing adversaries with a broader strategic space. Under specific settings of block timestamps, the UUM attack can exhibit similar or even equivalent characteristics to the RUM attack, providing an important perspective for us to deeply understand the feature of both attack methods.
Next, we will delve into the conditions under which the UUM attack downgrades into the RUM attack through Theorem 4.
Theorem 4 (UUM Downgrades to RUM).
The downgradation of UUM to RUM occurs if and only if $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor=1$.
Proof of Theorem 4.
The detailed proof of Theorem 4 can be found in Appendix E. ∎
IV-B State Space
Based on the analysis of the three phase strategy of the UUM attack, its state space can be divided into two categories: the deployment state and the attack state. The deployment state refers to the scenario where the UUM adversary waits for an appropriate opportunity to launch an attack. At this time, the blockchain topology and the timestamp of the latest block do not meet the attack conditions. The latest block is generated by either the adversary or honest participants, and the time difference between it and the previous main chain block is less than 9 seconds. The attack state means that the blockchain topology and the timestamp of the latest block comply with the attack conditions described in Theorem 2, that is, the latest block is generated by honest participants and the time difference between it and the previous main chain block is greater than or equal to 9 seconds.
Regarding state transitions, in the deployment state, if honest participants generate a block with a time difference greater than or equal to 9 seconds, the state changes to the attack state; otherwise, it remains in the deployment state. In the attack state, if honest participants continue to generate blocks with a time difference greater than or equal to 9 seconds, the attack state is maintained. If the adversary generates a block and strategically sets the timestamp, the state reverts to the deployment state. In other cases, the state also changes to the deployment state. The state transition process of UUM can be found in Figure 2-(b) and Table I.
IV-C Reward Analysis
By cleverly manipulating block timestamps and employing attack strategies, UUM adversaries attempt to obtain rewards exceeding their fair share, which inevitably impacts the interests of honest participants. Next, we conduct an in-depth analysis through Theorem 5 to examine the changes in the share of coin-base rewards between adversaries and honest participants under the UUM attack.
Prior to proving this theorem, we define some additional notations: $\mathbb{R}_{\mathcal{P}}^{RUM}$ denotes the absolute share of the coin-base reward for an RUM participant $\mathcal{P}$, and $\mathbb{E}\left[\mathbb{R}_{\mathcal{P}}^{RUM}\right]$ represents the expected relative share of the coin-base reward for an RUM participant $\mathcal{P}$.
Theorem 5 (UUM Adversary Rewards).
The expected relative share of coin-base rewards for UUM adversaries increases, while the absolute share remains unchanged. Conversely, both the expected relative and absolute shares of coin-base rewards for honest participants will decrease.
Proof of Theorem 5.
Based on the analysis of state transitions for UUM deployment and attack states presented in Section IV-B, we calculate the absolute and relative shares of coin-base rewards for both UUM adversaries and honest participants.
The reward analysis for all state transitions is analogous to that of selfish mining. Among these, a particular case warrants attention, corresponding to the transition from the Attack State to the Deployment State. In this case, the UUM adversary $\mathcal{P}_{\mathcal{A}}$ discovers the next valid block and publishes it, carefully setting the timestamp to ensure that the block’s difficulty exceeds that of an honest block. This manipulation causes the adversary’s block to be preferentially selected by other honest participants. Consequently, $\mathcal{P}_{\mathcal{A}}$ prevails in the fork competition and earns a coin-base reward. Conversely, the block generated by honest participants becomes an orphan block, triggering a transition from the Deployment State to the Attack State. This transition results in the recall of a coin-base reward that was prematurely awarded to honest participants. Therefore, the reward $\mathcal{P}_{\mathcal{H}}$ for honest participants corresponding to this state transition is 1-1.
The detailed proof of Theorem 5 can be found in Appendix F. ∎
The key mechanism lies in the irreversible shift of the steady-state probability:
  1. $\bullet$
    Persistent Existence of the Attack State. The deployment condition of the UUM attack ($\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\rfloor\geq 1$) is frequently triggered during the normal operation of the blockchain, causing the system to remain in the attack state for an extended period ($P(\text{Attack})\uparrow$).
  2. $\bullet$
    Self-Reinforcing Cycle. When the attack is successful, the adversary suppresses the growth of subsequent block difficulty to the lowest level through timestamp calibration (see Theorem 3), making the attack cost approach zero ($AC^{\text{UUM}}\approx 0$). This enables the adversary to maintain the attack state indefinitely, forming a positive feedback loop of attack $\to$ reward extraction $\to$ controllable difficulty risk $\to$ continuous attack.
  3. $\bullet$
    Zero-Sum Redistribution of Rewards. Each time the adversary succeeds, the absolute reward share of honest participants decreases by 1, while the relative share of the adversary grows linearly with the steady-state probability ($E[R_{P_{A}}^{\text{UUM}}]\propto P(\text{Attack})$). Since the difficulty adjustment cannot automatically correct this deviation, the reward distortion will be permanently embedded in the protocol’s economic model.
TABLE I: UUM state transition.
State Destination Transition Reward
$\mathcal{P}_{\mathcal{H}}$ $\mathcal{P}_{\mathcal{S}}$
Deployment Deployment $\mathcal{P}_{\mathcal{A}}~~any~t$ 0 11
Deployment Deployment $\mathcal{P}_{\mathcal{H}}~~t<9$ 11 0
Deployment Attack $\mathcal{P}_{\mathcal{H}}~~t\geq 9$ 11 0
Attack Attack $\mathcal{P}_{\mathcal{H}}~~t\geq 9$ 11 0
Attack Deployment $\mathcal{P}_{\mathcal{H}}~~t<9$ 11 0
Attack Deployment $\mathcal{P}_{\mathcal{A}}~~any~t$ 1-1 11
  • * The state transition represented by the gray row (the sixth row) is the fundamental cause of UUM adversaries gaining an advantage over RUM and honest participants.
V The Design of SUUM Attack
The SUUM attack expands the strategy space for adversaries beyond the RUM attack by removing the constraint in UUM that requires adversary to immediately release discovered blocks. Unlike UUM, where adversary has to release a block as soon as it is discovered, adversaries in SUUM have the freedom to not only carefully manipulate block timestamps but also to decide whether to withhold or release the generated blocks. This flexibility enhances the strategy space for SUUM adversaries. Compared to the UUM attack, the SUUM attack demonstrates stronger persistence and can generate higher excess profits for the adversary.
To achieve minimal risk control, SUUM adversaries need to more precisely select block timestamps during the attack process and optimize the timing and manner of their attack behaviors. By cleverly manipulating timestamps, adversaries can ensure the success of their attacks while minimizing the impact on the growth of blockchain difficulty, making their attack behaviors more concealed and sustainable.
The SUUM attack achieves persistent reward maximization through three core synergistic strategies embedded in its five phase framework, formally defined in Theorem 6-7 and visualized in Figure 3-(c). These strategies create a self-reinforcing loop by integrating timestamp manipulation, block withholding, and difficulty risk control:
  1. $\bullet$
    Timestamp Manipulation for Difficulty Dominance. During the Deployment Phase, once the adversary withholds the first private block $B_{1}^{p_{a}}$, timestamp manipulation becomes critical for difficulty advantage. By setting $t_{1}^{p_{a}}-t_{0}^{p_{a}}\in[1,9)$ (Theorem 6 Condition 1), the adversary ensures: $\mathcal{D}_{1}^{p_{a}}=\mathcal{D}_{0}^{p_{a}}+\left(1-\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right\rfloor\right)\cdot\left\lfloor\frac{\mathcal{D}_{0}^{p_{a}}}{2048}\right\rfloor$, where $\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right\rfloor=0$ (due to $t<9$), yielding a positive difficulty increment. In contrast, honest blocks with $t_{1}^{p_{h}}-t_{0}^{p_{h}}\geq 9$ incur a non-positive increment, ensuring $\mathcal{D}_{1}^{p_{a}}>\mathcal{D}_{1}^{p_{h}}$ and preferential chain selection (Figure 3-(c) transition from Attack State to Release Phase).
  2. $\bullet$
    Block Withholding for Reorganization Cascades. The withholding strategy extends to subsequent blocks ($i\geq 2$), where the adversary appends $B_{i}^{p_{a}}$ to the private chain until an honest block $B_{i}^{p_{h}}$ emerges. By controlling timestamp differences to satisfy $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}-(t_{i}^{p_{a}}-t_{i-1}^{p_{a}})}{9}\right\rfloor\geq 1$ (Theorem 6 Condition 2), each released private block surpasses the honest chain’s difficulty, enabling cascading reorganizations. As shown in Figure 3-(c), the state transition from Attack $i$ to Attack $i+1$ represents ongoing withholding, while Attack $i$ to Deployment triggers strategic release, maximizing fork opportunities (Figure 6-(b)).
  3. $\bullet$
    Difficulty Risk Control for Cost-free Persistence. To minimize difficulty fluctuations, the adversary restricts $t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\in[1,9)$ (Theorem 7 Condition 1), ensuring $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i}^{p_{a}}}{9}\right\rfloor=1$ and minimal risk (maximal difficulty deviation 0.21, Figure 6-(a)). The state space’s Downgrade State (Figure 3-(c)) acts as a safety mechanism, allowing fallback to UUM strategies when honest timestamp differences exceed 9 seconds, maintaining a higher steady-state attack probability than UUM (Figure 4) and attack cost $\mathbb{A}\mathbb{C}^{SUUM}\approx 0$ (Theorem 12).
V-A Method
The SUUM attack achieves persistent reward maximization through a synergistic five-phase framework integrating timestamp manipulation, strategic block withholding, and dynamic difficulty risk control. Unlike the attack methodologies of RUM and UUM, SUUM introduces a novel capability for adversaries to withhold mined blocks, form private chains, and strategically release them in response to network dynamics. This empowers adversaries to delay block publication and orchestrate chain reorganizations with precision.
Specifically, the SUUM attack achieves the goal of low risk and high returns through five phases: SUUM downgrade phase, SUUM deployment phase, SUUM block withholding phase, SUUM block release phase, and SUUM minimum risk control phase.
  1. $\bullet$
    Phase One: SUUM Downgrade Phase. If adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM observes that the difference between the timestamp $t_{1}^{p_{h}}$ of a newly generated block $\mathcal{B}_{1}^{p_{h}}$ by honest participant $\mathcal{P}_{\mathcal{H}}$ in the network and the timestamp $t_{0}^{p_{h}}$ of the previous main chain block $\mathcal{B}_{0}^{p_{h}}$ is $t_{1}^{p_{h}}-t_{0}^{p_{h}}\in[9,+\infty)$, i.e., $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor\in[1,+\infty)$, then SUUM downgrades to UUM. Note that when $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor=1$, SUUM also downgrades to RUM. The subsequent strategies of adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM are identical to those in UUM. If adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM first discovers a valid block, it proceeds to Phase Two.
  2. $\bullet$
    Phase Two: SUUM Deployment Phase. At this point, adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM first discovers a valid block $\mathcal{B}_{1}^{p_{a}}$. $\mathcal{P}_{\mathcal{A}}$ will withhold this block $\mathcal{B}_{1}^{p_{a}}$, forming a private chain visible only locally to himself, and then transitions to Phase Three.
  3. $\bullet$
    Phase Three: SUUM Withholding Block Phase. The adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM continues mining along the private chain. If $\mathcal{P}_{\mathcal{A}}$ discovers a new valid block $\mathcal{B}_{2}^{p_{a}}$, he appends this block to its local private chain and continues executing Phase Three. If, on the other hand, an honest participant $\mathcal{P}_{\mathcal{H}}$ discovers a new valid block $\mathcal{B}_{1}^{p_{h}}$, it transitions to Phase Four.
  4. $\bullet$
    Phase Four: SUUM Releasing Block Phase. At this juncture, the local private chain of adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM has a length of at least $1$, and honest participant $\mathcal{P}_{\mathcal{H}}$ discovers a new valid block $\mathcal{B}_{1}^{p_{h}}$. Adversary $\mathcal{P}_{\mathcal{A}}$ in SUUM releases a private block $\mathcal{B}_{1}^{p_{a}}$ and transitions to Phase Five. After the process of releasing the block is completed, if the length of $\mathcal{P}_{\mathcal{A}}$’s local private chain still remains at least $1$, he proceeds to Phase Three. If $\mathcal{P}_{\mathcal{A}}$ has no private blocks held in reserve, he transitions to Phase One.
  5. $\bullet$
    Phase Five: SUUM Minimum Risk Control Phase. $\mathcal{P}_{\mathcal{A}}$ carefully selects the timestamp $t_{1}^{p_{a}}$ for the block $\mathcal{B}_{1}^{p_{a}}$, aiming to achieve a higher difficulty $\mathcal{D}_{1}^{p_{a}}$ compared to the honest block’s difficulty $\mathcal{D}_{1}^{p_{h}}$ while minimizing the subsequent block difficulty growth rate. Upon completion of this step, $\mathcal{P}_{\mathcal{A}}$ proceeds with the subsequent steps of Phase Four.
Compared to the UUM attack, the most notable distinction of the SUUM attack lies in its allowance for adversaries to strategically withhold or release blocks under specific circumstances. This characteristic enables adversaries to more flexibly respond to various situations within the blockchain network, increasing the likelihood of attack success and potential rewards. To fully grasp the mechanism of the SUUM attack, we first clarify its success conditions in Theorem 6.
TABLE II: SUUM state transition.
State Destination Transition Reward
$\mathcal{P}_{\mathcal{H}}$ $\mathcal{P}_{\mathcal{S}}$
Deployment Attack 11 $\mathcal{P}_{\mathcal{A}}~~any~t$ 0 11
Deployment Deployment $\mathcal{P}_{\mathcal{H}}~~t<9$ 11 0
Deployment Downgrade $\mathcal{P}_{\mathcal{H}}~~t\geq 9$ 11 0
Downgrade Deployment $\mathcal{P}_{\mathcal{H}}~~t<9$ 11 0
Downgrade Downgrade $\mathcal{P}_{\mathcal{H}}~~t\geq 9$ 11 0
Downgrade Deployment $\mathcal{P}_{\mathcal{A}}~~any~t$ 1-1 11
Attack 11 Deployment $\mathcal{P}_{\mathcal{H}}~~any~t$ 0 0
Attack i1i-1, $i\geq 2$ Attack $i$ $\mathcal{P}_{\mathcal{A}}~~any~t$ 1-1 11
Attack $i+1$,
$i\geq 1$
Attack $i$ $\mathcal{P}_{\mathcal{H}}~~any~t$ 11 0
  • * The state transitions that lead to honest blocks becoming orphaned include the two gray rows (the sixth and eighth rows).
Theorem 6 (SUUM Successful Condition).
The SUUM attack is successful if and only if the following conditions are met:
  1. (1)
    For $i=1$, the conditions $\left[\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{a}}}{9}\right]=\left[\frac{t_{1}^{p_{h}}-t_{0}^{p_{a}}}{9}\right]\in[1,+\infty)$ and $t_{i}^{p_{a}}-t_{i-1}^{p_{a}}=t_{1}^{p_{a}}-t_{0}^{p_{a}}\in[1,900)$ must be satisfied.
  2. (2)
    For $i\geq 2$, the conditions $\left[\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}-\left(t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\right)}{9}\right]\in[1,+\infty)$ and $t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\in[1,900)$ must be satisfied.
Proof of Theorem 6.
The detailed proof of Theorem 6 can be found in Appendix G. ∎
After thoroughly discussing the success conditions of the SUUM attack, we further focus on the minimum risk control strategy. Due to its more complex attack pattern involving operations such as withholding and releasing blocks, risk control becomes particularly important in the SUUM attack. Similar to the UUM attack, the primary risk associated with the SUUM attack manifests in its impact on blockchain difficulty.
Next, we will elaborate on the specific conditions of the SUUM attack in terms of minimal risk control through Theorem 7, revealing the internal mechanisms and characteristics of the SUUM attack under risk control strategies.
Theorem 7 (SUUM Successful Condition with Minimal Risk).
The SUUM attack is successful with minimal risk if and only if the following conditions are met:
  1. (1)
    For $i=1$, the conditions $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{a}}}{9}\right\rfloor=\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{a}}}{9}\right\rfloor=1$ and $t_{i}^{p_{a}}-t_{i-1}^{p_{a}}=t_{1}^{p_{a}}-t_{0}^{p_{a}}\in[1,900)$ must be satisfied.
  2. (2)
    For $i\geq 2$, the conditions $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}-\left({t_{i}^{p_{a}}-t_{i-1}^{p_{a}}}\right)}{9}\right\rfloor=1$ and $t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\in[1,900)$ must be satisfied.
Proof of Theorem 7.
The detailed proof of Theorem 7 can be found in Appendix H. ∎
V-B State Space
Based on the analysis of the three-phase SUUM attack strategy in Section V-A, macroscopically, we categorize the state space of SUUM into three types: deployment state, downgraded state, and attack state. Microscopically, we further divide the attack state into withholding state and releasing state. Specifically, the deployment state captures the scenario where the SUUM adversary is waiting for an appropriate opportunity, the downgrade state captures the situation where SUUM is downgraded to UUM, and the attack state captures the execution of attacks by the SUUM adversary, namely, strategically withholding or releasing blocks. We show the state transition process of SUUM in Figure 2-(c) and Table II. The details of each state are outlined below:
  1. $\bullet$
    SUUM Deployment State. Three subsequent scenarios may arise: 1) The SUUM adversary continues to wait until a latest block is generated by an honest participant, with a timestamp difference of greater than or equal to 9 compared to the previous main chain block. Upon fulfillment of this condition, the Deployment State transitions to the Downgrade State. 2) If the SUUM adversary successfully generates the next valid block, they will withhold it and drive the Deployment State to transition to Attack State 1. 3) If the next block is generated by an honest participant and the timestamp difference is less than $9$ compared to the previous mainchain block, the state remains unchanged, remaining in the Deployment State.
  2. $\bullet$
    SUUM Downgrade State. Three subsequent scenarios may arise: 1) If the SUUM adversary successfully generates a valid block, he will strategically set the block’s timestamp and drive the downgraded state to transition back to the Deployment State. 2) If the next block is generated by an honest participant and its timestamp difference is less than $9$ compared to the previous block, this drives the Downgrade State to transition back to the Deployment State. 3) If the next block is generated by an honest participant and its timestamp difference is greater than or equal to $9$ compared to the previous block, the state remains unchanged, remaining in the Downgrade State.
  3. $\bullet$
    SUUM Attack State. Three subsequent scenarios may arise: 1) For Attack State $1$, if the next valid block is generated by an honest participant, the SUUM adversary immediately releases a withheld block with a strategically set timestamp and drives the Attack state to transition back to the Deployment State. 2) For Attack State $i$, $i\geq 1$, if the next block is generated by the SUUM adversary, he will withholds it and drive the Attack State to transition back to the Deployment State. 3) For Attack State $i+1$, $i\geq 1$, if the next block is generated by an honest participant, the SUUM adversary immediately releases a withheld block and strategically sets its timestamp. The Attack State transitions back to the Deployment State.
V-C Reward Analysis
After a detailed analysis of the mechanisms of the SUUM attack, including its success conditions and minimal risk success conditions, we focus our attention on the impact of the SUUM attack on the reward distribution pattern within blockchain networks. As we observed in our study of the UUM attack, any attack behavior has the potential to disrupt the originally fair reward distribution mechanism designed in blockchain systems, and the SUUM attack is no exception, with its impact being even more complex and far-reaching.
Theorem 8 (SUUM Adversary Rewards).
The expected relative share of the mainchain block coin-base reward for the SUUM adversary increases, while the absolute share remains unchanged. In contrast, both the expected relative share and the absolute share of the coin-base reward for honest participants will decrease.
Proof of Theorem 8.
Based on the analysis of state transitions in the Deployment State, Downgrade State, and Attack State of SUUM presented in Section V-B, we calculate the absolute and relative shares of coin-base rewards for both the SUUM adversary and honest participants.
The reward analysis for all state transitions is analogous to that of the UUM attack. Three special cases deserve attention, corresponding respectively to the transitions from the Downgrade state to the Deployment state, from the Attack $1$ state to the Deployment state, and from the Attack $i,~i\geq 1$ state to the Attack $i+1$ state.
In the first case, the SUUM adversary $\mathcal{P}_{\mathcal{A}}$, while in the Downgrade state, discovers the next valid block and publishes it. By carefully setting the timestamp, $\mathcal{P}_{\mathcal{A}}$ ensures that the block’s difficulty exceeds that of honest blocks, causing the adversary’s block to be preferentially selected by other honest participants. Consequently, $\mathcal{P}_{\mathcal{A}}$ wins the fork competition and earns a coin-base reward. The block generated by the honest participant coalition becomes an orphan block, prompting a transition to the Attack state and necessitating the recall of a pre-paid coin-base reward intended for honest participants. Therefore, the reward $\mathcal{P}_{\mathcal{H}}$ for honest participants corresponding to this state transition is 1-1.
In the second case, an honest participant discovers and publishes a valid block while in the Attack $1$ state. Immediately, the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ releases a withheld block, carefully setting its timestamp to ensure its difficulty exceeds that of the honest block. Ultimately, the selfish player wins the fork competition, rendering the honest block an orphan, and $\mathcal{P}_{\mathcal{A}}$ earns a coin-base reward (pre-paid during the transition from the Deployment state to the Attack $1$ state). The honest player receives no reward and triggers a transition from the Attack $1$ state to the Deployment state.
In the third case, when the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ is in the Attack $i,~i\geq 1$ state and discovers the next valid block, $\mathcal{P}_{\mathcal{A}}$ withholds it and transitions the current state to the next state, Attack $i+1$. Regardless of subsequent blockchain evolutions, by meticulously setting the timestamp, $\mathcal{P}_{\mathcal{A}}$ ensures that the withheld block ultimately becomes part of the main chain. Consequently, $\mathcal{P}_{\mathcal{A}}$ earns a coin-base reward while causing a corresponding loss of a coin-base reward for competing honest participants.
The detailed proof of Theorem 8 can be found in Appendix I. ∎
VI Comparison
Compared to the RUM attack, the UUM attack significantly expands the strategic space. The RUM attack emphasize risklessness, with relatively strict attack conditions, whereas the UUM attack allows adversaries to obtain higher potential rewards while bearing a certain level of risk. This change is primarily reflected in the relaxation of conditions during the deployment and execution stages of the UUM attack, enabling adversaries to initiate attacks under a wider range of network states. From the perspective of reward distribution, the UUM attack results in an increase in the expected relative share of adversaries and a decrease in the share of honest participants, disrupting the original reward balance based on fair computing power input.
The SUUM attack further deepens the strategic capabilities of adversaries based on the UUM attack. They introduce strategies for withholding and releasing blocks, enabling adversaries to more flexibly respond to dynamic changes in blockchain networks. In terms of success conditions and minimal risk success conditions, the SUUM attack involves more complex block timestamp relationships and phase judgments, reflecting the high complexity of their attack strategies. Similar to the UUM attack, the SUUM attack also alters the reward distribution pattern, benefiting adversaries while damaging honest participants. However, due to its more powerful strategic space, the SUUM attack has a more profound impact on blockchain networks. Next, we detailedly compare the advantages of these attack methods in terms of rewards in Theorem 9.
Theorem 9 (Reward Comparison of Uncle Maker-based Attacks and Honest Mining).
SUUM outperforms UUM, which in turn outperforms RUM, all of which surpass honest mining.
Proof of Theorem 9.
The detailed proof of Theorem 9 can be found in Appendix J. ∎
The RUM attack maintains a probabilistically negligible difficulty risk due to its constrained operation, which inherently limits both its attack surface and potential impact on blockchain difficulty adjustment. In contrast, the UUM attack’s relaxation of temporal constraints introduces measurable but bounded difficulty risk, primarily determined by the Markovian state transition probabilities post-attack. The SUUM attack exhibits the highest difficulty risk profile, as its compounded strategy of staggered block release and timestamp manipulation in both attack and downgrade states creates non-linear interactions with the difficulty adjustment algorithm. Next, we use Theorem 10 to elaborate on the differences in blockchain difficulty risk between these attack methods and honest mining.
Theorem 10 (Difficulty Risk Comparison of Uncle Maker-based Attacks and Honest Mining).
Compared to honest mining, the difficulty risks posed by RUM, UUM, and SUUM attacks on blockchain difficulty are as follows:
  1. (1)
    The increased difficulty risk associated with the RUM attack is denoted by $\sum\left({Attack\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy}\right)$, which represents the sum of the instances where a RUM adversary successfully attacks and drives the Attack State to transition to the Deployment State..
  2. (2)
    The increased difficulty risk associated with the UUM attack is denoted by $\sum\left({Attack\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy}\right)$, which represents the sum of the instances where a UUM adversary successfully attacks and drives the Attack State to transition to the Deployment State.
  3. (3)
    The increased difficulty risk associated with SUUM is denoted by $\sum\begin{pmatrix}{Downgrade\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy}\\ {+Attack~1\overset{\mathcal{P}_{\mathcal{H}}~~any~t}{\Rightarrow}Deploy}\end{pmatrix}$, which represents the sum of the instances of two types of state transitions corresponding to successful attacks by an SUUM adversary.
Proof of Theorem 10.
The detailed proof of Theorem 10 can be found in Appendix K. ∎
Through the preceding theoretical analysis of three types of Uncle Maker-based attacks, we understand that their essence lies in maliciously creating uncle blocks to induce forks in the blockchain and meticulously manipulating timestamps to prioritize the adoption of adversary blocks by other network participants. Consequently, the magnitude of the deliberate forking rate serves as a metric for assessing the severity of Uncle Maker-based attacks. Theorems 9 and 10 compare RUM, UUM, SUUM, and honest mining from the perspectives of reward and risk, respectively. Intuitively, as the maliciousness of RUM, UUM, and SUUM increases, their forking rates should also increase correspondingly. To streamline our notation, we use $\mathbb{P}_{S}^{A}$ to denote the steady-state probability of state $S$ when adversary $\mathcal{A}$ employs attack $A$, and ${\mathbb{F}\mathbb{R}}^{A}$ to represent the forking rate induced by attack $A$. In the following, we formally present a comparison of RUM, UUM, SUUM, and honest mining in terms of block forking rates.
Theorem 11 (Forking Rate Comparison of Uncle Maker-based Attacks and Honest Mining).
For an adversary $\mathcal{A}$ with the same computational power, the block forking rates induced by adopting RUM, UUM, SUUM, and honest mining satisfy the following relationship:
${\mathbb{F}\mathbb{R}}^{SUUM}>{\mathbb{F}\mathbb{R}}^{UUM}>{\mathbb{F}\mathbb{R}}^{RUM}>{\mathbb{F}\mathbb{R}}^{HM}.$ (1)
Specifically,
  1. (1)
    For honest mining, ${\mathbb{F}\mathbb{R}}^{HM}=0$.
  2. (2)
    For RUM, ${\mathbb{F}\mathbb{R}}^{RUM}=\mathbb{P}_{Attack}^{RUM}\cdot\mathcal{P}_{\mathcal{A}}(t<9)$, where $\mathcal{P}_{\mathcal{A}}(t<9)$ represents a scenario where an adversary $\mathcal{A}$ in RUM attack state finds the next block, and the difference in timestamps between it and its parent block is less than $9$.
  3. (3)
    For UUM, ${\mathbb{F}\mathbb{R}}^{UUM}=\mathbb{P}_{Attack}^{UUM}\cdot\mathcal{P}_{\mathcal{A}}(any~t)$, which indicates a scenario where an adversary $\mathcal{A}$ in the UUM attack state finds the next block, regardless of the difference in timestamps between it and its parent block.
  4. (4)
    For SUUM, ${\mathbb{F}\mathbb{R}}^{SUUM}=\mathbb{P}_{Attack}^{SUUM}+\mathbb{P}_{Downgrade}^{SUUM}\cdot\mathcal{P}_{\mathcal{A}}(any~t)$, which represents the sum of the steady-state probability of being in the attack state and the probability of an adversary in the downgrade state finding the next block, regardless of the difference in timestamps between it and its parent block.
Proof of Theorem 11.
The detailed proof of Theorem 11 can be found in Appendix L. ∎
Next, we focus our attention on the attack cost associated with three types of attacks. It is noteworthy that we characterize the attack cost as the reduction in the adversary’s block generation probability resulting from an increase in block difficulty. Prior to the detailed analysis, we first introduce some additional notations. We denote $\mu_{\mathcal{A}}$ as the proportion of power controlled by adversary $\mathcal{A}$, max_targetmax\_target as the maximum target value, and $\mathbb{A}\mathbb{C}$ as the attack cost.
Subsequently, in Theorem 12, we formally present the attack cost for these three types of attacks.
Theorem 12 (Cost Comparison of Uncle Maker-based Attacks).
For an adversary $\mathcal{A}$ possessing the same relative power, the attack cost of RUM, UUM, and SUUM satisfy the following relationship:
${\mathbb{A}\mathbb{C}}^{SUUM}={\mathbb{A}\mathbb{C}}^{UUM}\approx{\mathbb{A}\mathbb{C}}^{RUM}=0.$ (2)
Specifically,
  1. (1)
    For RUM, ${\mathbb{A}\mathbb{C}}^{RUM}=0$.
  2. (2)
    For UUM or SUUM, ${\mathbb{A}\mathbb{C}}^{UUM}={\mathbb{A}\mathbb{C}}^{SUUM}=\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}_{0}^{{p}_{{h}}}-\mathcal{D}_{1}^{{p}_{{h}}}}{max\_target}\leq\mu_{\mathcal{A}}\cdot\frac{1}{2^{216.35}}\approx 0.$
Proof.
The detailed proof of Theorem 12 can be found in Appendix M. ∎
This result once again implies that adversaries can relatively easily carry out UUM and SUUM attacks with almost no cost risk, thereby further highlighting the severe threats posed by these two attack strategies to the security of timestamp-based Nakamoto-style blockchains. Since nearly cost-free attacks may prompt more adversaries to attempt to exploit these vulnerabilities, undermining the fairness and stability of the blockchain.
Figure 3: Steady-state Probability. This figure illustrates the steady-state probabilities of three different attack strategies as a function of the adversary’s relative power $\alpha$. In the context of RUM and UUM, the steady-state probability of the Attack State serves as a critical metric for quantifying the detrimental impact of attacks. Specifically, a higher steady-state probability of the Attack State directly correlates with a more severe level of harm inflicted on the system. Conversely, for SUUM, the overall harm magnitude is jointly determined by the steady-state probabilities of three distinct states. Note that we only show the total probability of all Attack states within the SUUM model.
VII Simulated Estimation
To empirically validate the theoretical analysis of the UUM and SUUM attacks, we conducted extensive simulations under various adversarial power ratios. These simulations aimed to quantify the steady-state probabilities, reward distributions, difficulty risks, and forking rates associated with each attack strategy. By comparing these metrics with honest mining and the baseline RUM attack, we demonstrate the enhanced profitability and persistence of the proposed advanced variants. The simulation setup and results are detailed in the following subsections.
Similar to [26, 32], to reduce the waste of computational resources, we did not conduct experiments on the real Ethereum 1.x system. To achieve the same goal, we implemented a discrete-event simulator to model the Ethereum 1.x blockchain protocol, incorporating the fork selection rules and difficulty adjustment mechanisms described in Section III. The simulator tracks block generation, timestamp manipulation, and adversarial strategies under controlled conditions. Key parameters include:
  1. $\bullet$
    Adversarial power ratio $\bm{\alpha}$: Varied from 0 to 0.5 in increments of 0.05.
  2. $\bullet$
    Block generation: Modeled as a Poisson process with a 13-second average block time, consistent with Ethereum 1.x.
  3. $\bullet$
    Timestamp constraints: Enforced Ethereum 1.x style blockchain validity rules.
  4. $\bullet$
    Difficulty adjustment: Calculated dynamically based on block timestamps and parent difficulties.
Each simulation ran for 1M blocks to ensure convergence to steady-state behavior, with results averaged over 10K trials to minimize variance. The following subsections present the findings for each evaluated metric.
VII-A Estimate the Steady-state Probability
To quantitatively analyze the effectiveness of the proposed attack strategies, we estimate the steady-state probabilities of the RUM, UUM, and SUUM attacks under varying adversarial power ratios. The steady-state probability reflects the long-term likelihood of the system being in an attack state, which directly correlates with the adversary’s ability to sustain the attack and maximize rewards.
As shown in Figure 3, the steady-state probability of the deployment state increases with the adversary’s power for RUM and UUM strategies. However, SUUM exhibits the highest steady-state probability of the deployment state across higher power levels, followed by UUM and RUM. This is because SUUM’s ability to withhold and strategically release blocks expands its attack opportunities, while UUM’s relaxation of risk-free constraints allows more frequent transitions to the attack state compared to RUM.
The results validate that SUUM’s design achieves superior persistence in maintaining the attack state, enabling adversaries to exert sustained influence on the blockchain. This aligns with Theorem 7, where SUUM’s reward advantage stems from its higher steady-state probability of attack execution. The findings underscore the need for countermeasures to mitigate such persistent attacks, as discussed in Section VIII.
(a) $RR_{Attack}$
(b) $RR_{Honest}$
Figure 4: Comparison of Relative Rewards under Different Mining Strategies. (a) Comparison of Adversary Relative Rewards under Different Mining Strategies. This figure compares the relative reward gains of adversaries under different mining strategies, illustrating the reward disparities among the SUUM, UUM, and RUM attack strategies compared to honest mining. The results demonstrate that SUUM yields the highest rewards, significantly outperforming UUM and RUM, while all three attack strategies surpass honest mining in profitability. This outcome confirms that adversaries can obtain excess rewards by manipulating timestamps and strategically withholding blocks. (b) Comparison of Honest Participant Relative Rewards under Different Mining Strategies. This figure presents a comparative analysis of honest participants’ relative rewards under different mining strategies. Notably, SUUM exhibits the most severe reward suppression effect, followed by UUM and RUM, with all three attack strategies significantly undercutting the baseline rewards achievable through honest mining.
VII-B Estimate the Relative Reward
To evaluate the economic impact of the proposed attacks, we analyze the relative rewards obtained by adversaries and honest participants under RUM, UUM, SUUM, and honest mining strategies. The relative reward measures the proportion of total block rewards captured by each party, reflecting the fairness and security of the blockchain’s incentive mechanism.
The relative rewards for adversaries ($RR_{Attack}$) and honest participants ($RR_{Honest}$) are calculated by dividing their earned rewards by the total system rewards. As shown in Figure 4, SUUM consistently yields the highest relative rewards for adversaries, surpassing UUM and RUM across all power levels. For instance, at $\alpha=0.25$, SUUM enables adversaries to capture 33.30% of the rewards, while UUM and RUM achieve 28.41% and 26.12%, respectively. This aligns with Theorem 9, where SUUM’s block withholding and timestamp manipulation synergistically amplify rewards. Honest mining, as expected, adheres to the fair share, serving as the baseline.
Figure 4 demonstrates the corresponding degradation in honest participants’ rewards under adversarial strategies. SUUM inflicts the most severe suppression, reducing honest rewards by up to 11.85% compared to honest mining at $\alpha=0.3$. UUM and RUM exhibit intermediate effects, with honest rewards declining linearly as $\alpha$ increases. This inverse relationship between adversarial and honest rewards confirms the zero-sum nature of reward redistribution in these attacks.
The death spiral effect of SUUM, emerges from a self-reinforcing cycle where adversarial rewards come at the direct expense of honest participants, driving a catastrophic breakdown of the protocol’s economic model.
VII-C Estimate the Minimal Difficulty Risk
A critical aspect of Uncle Maker attacks is their impact on blockchain difficulty, which influences long-term network stability. Here, we evaluate the minimal difficulty risk-the least additional risk imposed on the blockchain’s difficulty adjustment mechanism by each attack strategy. This metric reflects how subtly adversaries can execute attacks without destabilizing the network.
The minimal difficulty risk is quantified as $MR=\mathcal{D}_{Attack}-\mathcal{D}_{Honest},$ where $\mathcal{D}_{Attack}$ and $\mathcal{D}_{Honest}$ represent the difficulty of adversarial and honest blocks at the same highest height. Results are averaged across 10K trials to ensure statistical robustness.
Figure 5 reveals distinct risk profiles for each strategy. Honest Mining maintains a baseline risk of zero, as no difficulty manipulation occurs. RUM introduces negligible risk, as its strict timestamp constraints (Theorem 3) limit difficulty fluctuations. UUM exhibits marginally higher risk due to relaxed initiation conditions (Theorem 1), allowing occasional difficulty spikes. SUUM poses the highest risk, as its block withholding strategy (Section V-A) disrupts difficulty adjustment continuity. We further heuristically find three observations:
  1. $\bullet$
    Low Difficulty Risk Attacks: All strategies induce sub-0.21 difficulty risk, confirming that Uncle Maker attacks are low difficulty risk.
  2. $\bullet$
    Trade-off with Profitability: SUUM’s higher risk aligns with its superior rewards (Section VII), demonstrating a risk-reward balance.
  3. $\bullet$
    Network Stability: Even SUUM’s maximal difficulty risk remains manageable (e.g., <0.21<0.21 at $\alpha=0.37$), explaining why such attacks could persist undetected in practice.
Crucially, SUUM’s minimal difficulty risk allows this death spiral to proceed undetected, as the attack remains cost-free (Theorem 12). By calibrating timestamps to 1-second granularity (Section 5.1), adversaries avoid triggering difficulty spikes that would otherwise penalize their mining efficiency. This stealth enables sustained reward extraction, as seen in Figure 5-(b). The result is a low-risk, high-reward environment that incentivizes even moderate power miners to defect, amplifying the spiral.
VII-D Estimate the Forking Rate
SUUM’s elevated forking rate directly correlates with honest reward suppression in Figure 5-(b). Each fork invalidates honest blocks, reducing their effective hash power contribution and further discouraging participation. This creates a vicious cycle: higher forking $\to$ fewer honest blocks $\to$ lower incentives $\to$ more miners abandon honesty, as reflected in the steep decline of honest relative rewards for SUUM compared to UUM/RUM.
We simulate RUM, UUM, SUUM, and honest mining over 1M blocks, measuring the forking rate ($FR$), where the results are averaged across 10K trials, with adversarial power $\alpha$ varying from 0 to 0.5. We note that $FR$ is defined as the number of intentionally induced forks divided by the total number of blocks mined in the system.
Figure 5 highlights stark contrasts in forking behavior. Honest Mining maintains a 0% forking rate, as all participants adhere to the canonical chain. RUM triggers occasional forks, occurring only when adversaries mine blocks with timestamps $t<9$ (Theorem 11). UUM exhibits higher fork rates, as its relaxed constraints (Section IV-A) permit more frequent adversarial interventions. SUUM maximizes forks by combining withheld block releases (Section V-A) and timestamp manipulation, amplifying chain reorganizations.
The death spiral effect underscores SUUM’s unique threat: unlike prior attacks, it does not merely increase adversarial rewards but systematically erodes the incentive for honest participation, leading to irreversible protocol collapse. The combination of cost-free persistence (Theorem 12), recursive timestamp manipulation, and cascading reorganizations creates a feedback loop that accelerates as more miners defect. This highlights the urgent need for mitigations that break this cycle (Section 8), to restore honest miners’ incentives and prevent systemic failure.
(a) MRMR
(b) FRFR
Figure 5: Comparison of Minimal Difficulty Risk and Forking Rates under Different Attack Strategies. (a) Comparison of Minimal Difficulty Risk under Different Attack Strategies. This figure presents a comparative assessment of the minimal difficulty risk levels associated with different attack strategies. Honest mining maintains a baseline risk level of zero, while RUM introduces only minimal difficulty escalation risk. In contrast, UUM exhibits marginally higher risk due to its less restrictive attack conditions, and SUUM demonstrates the most significant risk elevation attributable to its block withholding mechanism. (b) Comparison of Forking Rates under Different Mining Strategies. This figure presents a comparative analysis of forking rates under different attack strategies, illustrating the relationship between the adversary’s relative power and the resultant forking rate. The results demonstrate that honest mining maintains a zero forking rate due to strict protocol compliance, while the RUM, UUM, and SUUM attacks exhibit progressively increasing fork probabilities.
VII-E Empirical Validation on Real Blockchain Networks
Previous sections have validated the theoretical effectiveness of SUUM/UUM attacks through discrete-event simulations (Sections VII.A–VII.D). However, the real-world existence of such attacks requires corroboration from actual blockchain data. This section leverages on-chain data from three ETH 1.x-style blockchains, Ethereum (ETH, block heights 15505647–15535776, 30,130 mainchain blocks), Ethereum Classic (ETC, block heights 23232147–23242146, 10,000 mainchain blocks), and Ethereum PoW (block heights 22905613–22915612, 10,000 mainchain blocks), to address three key questions via two experiments: 1) Whether the distribution of timestamp differences aligns with the attack window characteristics of SUUM/UUM; 2) Whether top mining pools engage in systematic timestamp manipulation; 3) Whether certain pools deliberately avoid timestamp differences divisible by 9. Ultimately, these experiments confirm the real-world pervasiveness of SUUM/UUM attacks.
Timestamp Difference in Mainchain Blocks. We counted the number of mainchain blocks corresponding to timestamp differences from parent blocks across the three blockchains, with visualizations presented in Fig. 6a. The statistical results indicate that the proportion of mainchain blocks with timestamp differences divisible by 9 is significantly lower than the expected normal trend. This corroborates that some mining pools are still maliciously manipulating timestamps to gain disproportionate advantages. Furthermore, we observed that the decrease in the number of blocks with timestamp difference $t$ exactly matches the increase in the number of blocks with difference t1t-1. This pattern perfectly corresponds to the difficulty risk control mechanism discussed in previous sections.
Four Timestamp Manipulation Pools. For each of the three blockchains, we sorted mining pools in descending order of total blocks mined and selected the top 10 pools. Among these 30 top pools, we identified four pools that actually implement timestamp manipulation: Ethereum 1.x (0x829bd8…), Ethereum PoW (0x9205c2…), and Ethereum Classic (0x406177… and 0x35aa26…), with their timestamp difference distributions visualized in Fig. 6b. Notably, none of their timestamp difference distributions include values divisible by 9. When a mainchain block would otherwise have a timestamp difference divisible by 9, these pools maliciously adjust it to the original value minus 1. This minimizes difficulty risk while securing a dominant advantage in fork competitions, aligning perfectly with the proposed SUUM/SUM attack strategies.
(a) Three ETH 1.x Blockchains
(b) Four Malicious Pools
Figure 6: Circumstantial Evidence for Attack. (a) This figure presents the distribution of mainchain blocks across three ETH 1.x-style blockchains, Ethereum 1.x, Ethereum PoW, and Ethereum Classic with respect to their timestamp differences from parent blocks. Blocks corresponding to timestamp differences of 9, 18, and 27 seconds are significantly fewer than those in adjacent intervals, while blocks with differences of 8, 17, and 26 seconds exhibit a marked surge. This serves as direct evidence of mining pools’ deliberate manipulation to avoid timestamp differences divisible by 9. (b) This figure focuses on four malicious mining pools, presenting the distribution of timestamp differences among mainchain blocks mined by these pools. All four pools exhibit zero blocks with timestamp differences of 9, 18, and 27 seconds, while blocks corresponding to differences of 8, 17, and 26 seconds show a significant surge. This distribution characterized by the avoidance of multiples of 9 and concentration on values equal to multiples of 9 minus 1 serves as direct evidence of the strategy in SUUM attacks to minimize difficulty risks and maximize competitive advantages in fork contests. It confirms that these pools are systematically executing timestamp manipulation attacks.
VIII Discussion
We now analyze the essence of the timestamp manipulation and propose some mitigation measures to effectively prevent it and its advanced variants. These measures mainly focus on adjusting the consensus mechanism and incentive mechanism of the blockchain, aiming to reduce the profit space of adversaries and enhance the security and fairness of the timestamp-based Nakamoto-style blockchains. The details can be found in Appendix N
IX Conclusion
This study demonstrates that timestamp-based Nakamoto-style blockchains, particularly Ethereum 1.x derivatives, face existential threats from the SUUM attack. Unlike transient adversarial strategies, SUUM inflicts permanent systemic damage by irreversibly distorting the protocol’s incentive structure through three synergistic mechanisms: 1) precision timestamp manipulation to inflate adversarial difficulty advantages, 2) strategic block withholding to amplify chain reorganizations, and 3) granular difficulty risk control to ensure cost-free sustainability. These mechanisms create a self-reinforcing cycle: adversarial rewards scale super-linearly, while honest participants face diminishing returns, ultimately incentivizing rational miners to defect and accelerating protocol collapse.
Our simulations validate SUUM’s dominance over prior attacks. Crucially, SUUM bypasses traditional constraints like hash power thresholds by exploiting timestamp calibration, maintaining minimal difficulty escalation despite sustained exploitation. This asymmetric advantage stems from the protocol’s inability to self-correct reward distortions, leading to a death spiral where adversarial coalitions dominate and honest participation becomes economically untenable. We have exposed four mining pools that conduct timestamp manipulation attacks in three major Ethereum 1.x-based blockchain systems.
Appendix A Related Work
We now embark on a discussion concerning the related work of this paper, primarily focusing on three aspects: attacks targeting vulnerabilities in incentive models and consensus protocols of Nakamoto-like and Ethereum 2.x-like blockchains.
A-A Attacks on Incentive Model and Consensus Protocol of Nakamoto-like Blockchains
Withholding Attack. Withholding attacks are prevalent malicious behaviors in Nakamoto-like blockchains [25, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Most consensus-layer attacks (e.g., selfish mining, stubborn mining) rely on block withholding: adversaries withhold blocks to control release timing or continue mining on lagging selfish forks [34, 27], undermining network fairness, wasting honest nodes’ computing resources, and reducing network security and reliability.
Bribing Attack. Bribing attacks pose severe threats to the security and stability of Nakamoto-like blockchains [42, 43, 44, 45, 46, 47, 48]. First proposed by Bonneau, these attacks include multiple mechanisms (e.g., inbound/outbound bribery via smart contracts). Their motives range from double-spending and consensus disruption to censorship and HTLC sabotage [49, 50], significantly disrupting blockchain operations.
DoS attack. Denial of Service (DoS) attacks disrupt system availability by consuming network resources [51, 52, 53, 54]. At the network layer, adversaries often launch traffic flooding attacks (e.g., fake requests) to exhaust bandwidth, hindering legitimate transaction processing, increasing confirmation latency, and weakening blockchain decentralization.
Timestamp Attack. Timestamp manipulation Attacks exploit configurable block timestamps in Nakamoto-like blockchains [14], interfering with block generation/confirmation or manipulating difficulty adjustments. Research on timestamp attacks at the consensus layer remains limited: while Stretch identified timestamp vulnerabilities in Geth (e.g., reducing honest mining difficulty), the impact on consensus mechanisms was not explored. Note that the UUM attack proposed in this paper belongs to timestamp attacks, and SUUM combines withholding attacks with timestamp attacks.
A-B Attacks on Incentive Model and Consensus Protocol of Ethereum 2.x-like Blockchains
Withholding Attack. In Ethereum 2.x-like PoS blockchains, withholding attacks threaten network stability [55, 56, 57, 58]. Adversaries withhold blocks to block propagation and release them opportunistically, exploiting staking/voting mechanisms to disrupt block confirmation, affect block height/forks, and undermine network consistency.
Nothing-at-stake Attack. Nothing-at-stake attacks severely threaten the activity and performance of Ethereum 2.x-like PoS blockchains [59, 60, 61]. Leveraging PoS staking/validation flaws, adversaries participate in multiple forks without risk, causing fork chaos, hindering consensus unification, and reducing transaction processing efficiency.
Long-range Attack. Long-range attacks are potential threats to Ethereum 2.x-like PoS blockchains [62, 60, 63]. Adversaries secretly construct alternative chains from historical blockchain points (with low resource consumption), undermining data consistency (e.g., altering past transactions) and eroding the trust foundation of PoS blockchains.
DoS Attack. DoS attacks on Ethereum 2.x-like PoS blockchains take multiple forms [64, 65, 66]. Network-layer attacks involve traffic flooding, while application-layer attacks target smart contracts (e.g., infinite loops) and Txpool (e.g., low-gas high-complexity transactions), causing resource exhaustion or congestion.
Reorg Attack. Reorg attacks threaten the security and stability of Ethereum 2.x-like PoS blockchains [55, 67]. Adversaries increase the proportion of Byzantine validator blocks via controlled nodes (including Short/Long Reorg Attacks), replacing the main chain, making transaction states uncertain, and weakening user trust.
Appendix B Proof of Theorem 1
The initiation condition for the UUM attack is $\mathcal{D}_{1}^{{p}_{{h}}}\leq\mathcal{D}_{0}^{{p}_{{h}}}$. Based on the difficulty calculation formula, we have:
$\mathcal{D}_{0}^{{p}_{{h}}}+{m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor,-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{{p}_{{h}}}}{2048}\right\rfloor$ $\leq\mathcal{D}_{0}^{{p}_{{h}}}$ (3)
${m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor,-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{{p}_{{h}}}}{2048}\right\rfloor$ $\leq 0$
$\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor$ $\geq 1.$
Hence, we have $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$.
Appendix C Proof of Theorem 2
Firstly, for the adversary’s block $\mathcal{B}_{{i}}^{{p}_{{a}}}$ to be valid, it is necessary for block $\mathcal{B}_{{1}}^{{p}_{{a}}}$ to be generated after its parent block $\mathcal{B}_{{0}}^{{p}_{{a}}}$. Consequently, the timestamp ${t}_{1}^{{p}_{{a}}}$ of block $\mathcal{B}_{1}^{{p}_{{a}}}$ should be greater than the timestamp ${t}_{0}^{{p}_{{a}}}$ of its parent block $\mathcal{B}_{{0}}^{{p}_{{a}}}$, i.e.:
${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}\geq 1.$ (4)
Secondly, only when the difficulty $\mathcal{D}_{1}^{{p}_{{a}}}$ of the adversary’s block $\mathcal{B}_{1}^{{p}_{{a}}}$ is greater than the difficulty $\mathcal{D}_{1}^{{p}_{{h}}}$ of the honest block $\mathcal{B}_{0}^{{p}_{{h}}}$ will other honest participants choose the adversary’s block. At this point, we have: $\mathcal{D}_{1}^{{p}_{{a}}}>\mathcal{D}_{1}^{{p}_{{h}}}$. Based on the difficulty calculation formula, we can derive the following inequality:
$\mathcal{D}_{0}^{{p}_{{a}}}+{m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor,-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{{p}_{{a}}}}{2048}\right\rfloor$ (5)
$>\mathcal{D}_{0}^{{p}_{{h}}}+{m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{h}}}}{9}\right\rfloor,-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{{p}_{{h}}}}{2048}\right\rfloor.$
Since $\mathcal{D}_{0}^{{p}_{{h}}}=\mathcal{D}_{0}^{{p}_{{a}}}$ and ${t}_{0}^{{p}_{{h}}}={t}_{0}^{{p}_{{a}}}$, we therefore have:
${m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor,-99}\right\}$ (6)
$>{m}{a}{x}\left\{{1-\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor,-99}\right\}.$
The above inequality can be transformed into the following system of inequalities:
$\left\{\begin{matrix}{1-\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor>1-\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor}&(1)\\ {1-\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor>-99}&(2)\end{matrix}\right.$. (7)
Based on Inequality (1), we have:
$\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor-\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor>0.$ (8)
Further rearranging the above inequality, we can derive:
$\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor>0.$ (9)
Since $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor$ can only take integer values, i.e., $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\in{N}$, we therefore have:
$\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\geq 1.$ (10)
Based on the inequality above, we can get $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$.
Based on the Inequality (10), we can derive:
${t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}\geq 9.$ (11)
Based on the Inequality (2), we can derive:
$\begin{matrix}{\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor<100}\end{matrix}$ (12)
Since $\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor$ can only take integer values, i.e., $\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor\in{N}$, we therefore have: $\left\lfloor\frac{{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}}{9}\right\rfloor\leq 99$. Based on the aforementioned inequalities, we have:
${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}<900.$ (13)
Combining Inequality (4) and Inequality (13), we obtain: $1\leq{t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}<900$, which implies that ${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}\in[1,900)$.
Appendix D Proof of Theorem 3
According to Theorem 2, the conditions for a successful UUM attack are: $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$ and ${t}_{1}^{{p}_{{a}}}-{t}_{0}^{{p}_{{a}}}\in[1,900)$. To minimize the risk associated with the UUM attack, which corresponds to minimizing the block difficulty growth rate, Theorem 2 suggests that we take the minimum value within the condition $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor\in\left[1,+{\infty}\right)$, i.e., $\left\lfloor\frac{{t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}}{9}\right\rfloor=1$. This implies that ${t}_{1}^{{p}_{{h}}}-{t}_{1}^{{p}_{{a}}}\in[9,18)$. By integrating these two conditions, we can prove the theorem.
Appendix E Proof of Theorem 4
The condition for the RUM attack is that mining on top of the former block incurs no additional risk, i.e., $\mathcal{D}_{1}^{p_{h}}=\mathcal{D}_{0}^{p_{h}}$. Based on the difficulty calculation formula, we have:
$\mathcal{D}_{0}^{p_{h}}+max\left\{{1-\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor,-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{p_{h}}}{2048}\right\rfloor$ $=\mathcal{D}_{0}^{p_{h}}$ (14)
$max\left\{{1-\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor,-99}\right\}$ $=0$
$\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor$ $=1.$
Appendix F Proof of Theorem 5
Based on the analysis of state transitions for UUM deployment and attack states presented in Section IV, we calculate the absolute and relative shares of coin-base rewards for both UUM adversaries and honest participants.
The absolute share of coin-base rewards for UUM adversary $\mathcal{P}_{\mathcal{A}}$ is:
$\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{UUM}$ $=\mathbb{P}\left({Deploy}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$ (15)
$=\left({\mathbb{P}\left({Deploy}\right)+\mathbb{P}\left({Attack}\right)}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathbb{P}_{\mathcal{A}}\left({any~t}\right).$
Therefore, the absolute share of coin-base rewards for UUM adversary $\mathcal{P}_{\mathcal{A}}$ remains unchanged.
The absolute share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ is:
$\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{UUM}$ $=\mathbb{P}\left({Deploy}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({t<9}\right)+\mathbb{P}\left({Deploy}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)$ (16)
$+\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)+\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t<9}\right)$
$-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathbb{P}\left({Deploy}\right)\cdot\left({\mathcal{P}_{\mathcal{H}}\left({t<9}\right)+\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)}\right)$
$+\mathbb{P}\left({Attack}\right)\cdot\left({\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)+\mathbb{P}_{\mathcal{H}}\left({t<9}\right)}\right)$
$-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathbb{P}\left({Deploy}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)+\mathbb{P}\left({Attack}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\left({\mathbb{P}\left({Deploy}\right)+\mathbb{P}\left({Attack}\right)}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathcal{P}_{\mathcal{H}}\left({any~t}\right)-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$<\mathcal{P}_{\mathcal{H}}\left({any~t}\right).$
Therefore, the absolute share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ decreases.
The expected relative share of coin-base rewards for UUM adversary $\mathcal{P}_{\mathcal{A}}$ is:
$\mathbb{E}\left[\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{UUM}\right]$ $=\frac{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{UUM}}{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{UUM}+\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{UUM}}$ (17)
$=\frac{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}{\begin{pmatrix}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}\\ {-P\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}\end{pmatrix}}$
$>\frac{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}$
$=\mathbb{P}_{\mathcal{A}}\left({any~t}\right).$
Therefore, the expected relative share of coin-base rewards for UUM adversary $\mathcal{P}_{\mathcal{A}}$ increases.
The expected relative share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ is:
$\mathbb{E}\left[\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{UUM}\right]$ $=\frac{\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{UUM}}{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{UUM}+\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{UUM}}$ (18)
$=\frac{\mathcal{P}_{\mathcal{H}}\left({any~t}\right)-\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}{\begin{pmatrix}\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)\\ -\mathbb{P}\left({Attack}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)\end{pmatrix}}$
$<\frac{\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}$
$=\mathcal{P}_{\mathcal{H}}\left({any~t}\right).$
Therefore, the expected relative share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ decreases.
Appendix G Proof of Theorem 6
The timestamps of the honest block $\mathcal{B}_{i}^{p_{h}}$ and the SUUM adversary’s block $\mathcal{B}_{i}^{p_{a}}$ at the same height $i$ are denoted as $t_{i}^{p_{h}}$ and $t_{i}^{p_{a}}$, respectively, with corresponding difficulties $\mathcal{D}_{i}^{p_{h}}$ and $\mathcal{D}_{i}^{p_{a}}$. For the RUM attack to be successful, the difficulty $\mathcal{D}_{i}^{p_{a}}$ of the SUUM adversary’s block $\mathcal{B}_{i}^{p_{a}}$ at the same height should be greater than the difficulty $\mathcal{D}_{i}^{p_{h}}$ of the honest block $\mathcal{B}_{i}^{p_{h}}$. According to the difficulty calculation formula, we have:
$\mathcal{D}_{i}^{p_{h}}=\mathcal{D}_{i-1}^{p_{h}}+max\left\{1-\left[\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}}{9}\right],-99\right\}\cdot\left\lfloor\frac{\mathcal{D}_{i-1}^{p_{h}}}{2048}\right\rfloor$ (19)
and
$\mathcal{D}_{i}^{p_{a}}=\mathcal{D}_{i-1}^{p_{a}}+max\left\{1-\left[\frac{t_{i}^{p_{a}}-t_{i-1}^{p_{a}}}{9}\right],-99\right\}\cdot\left\lfloor\frac{\mathcal{D}_{i-1}^{p_{a}}}{2048}\right\rfloor.$ (20)
Given that $t_{0}^{p_{h}}=t_{0}^{p_{a}}$ and $\mathcal{D}_{0}^{p_{h}}=\mathcal{D}_{0}^{p_{a}}$ (representing the same block), for $i=1,~2~,\ldots,~n$, it is necessary to ensure that $\mathcal{D}_{i}^{p_{a}}>\mathcal{D}_{i}^{p_{h}}$ holds true consistently.
  1. (1)
    For $i=1$, we need to ensure that the following condition holds:
    $\mathcal{D}_{i}^{p_{a}}>\mathcal{D}_{i}^{p_{h}}\Rightarrow\mathcal{D}_{1}^{p_{a}}>\mathcal{D}_{1}^{p_{h}}.$ (21)
    According to the difficulty calculation formula, we have:
    $\mathcal{D}_{0}^{p_{a}}+max\left\{{1-\left[\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right],-99}\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{p_{a}}}{2048}\right\rfloor$ (22)
    $>\mathcal{D}_{0}^{p_{h}}+max\left\{1-\left[\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right],-99\right\}\cdot\left\lfloor\frac{\mathcal{D}_{0}^{p_{h}}}{2048}\right\rfloor$
    $\Rightarrow max\left\{{1-\left[\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right],-99}\right\}$
    $>max\left\{1-\left[\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right],-99\right\}.$
    By rearranging this inequality, we obtain the following system of inequalities:
    $\left\{\begin{matrix}{1-\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right\rfloor>1-\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{a}}}{9}\right\rfloor}&(1)\\ {1-\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right\rfloor>-99}&(2)\end{matrix}\right.$. (23)
    According to inequality (1), we have:
    $\begin{aligned} t_{1}^{p_{h}}-t_{0}^{p_{h}}-\left({t_{1}^{p_{a}}-t_{0}^{p_{a}}}\right)\geq 9\\ \Rightarrow t_{1}^{p_{h}}-t_{1}^{p_{a}}\geq 9\end{aligned}.$ (24)
    Thus, we have:
    $\left\lfloor\frac{t_{1}^{p_{h}}-t_{1}^{p_{a}}}{9}\right\rfloor\in[1,+\infty).$ (25)
    According to inequality (2), we have:
    $\left\lfloor\frac{t_{1}^{p_{a}}-t_{0}^{p_{a}}}{9}\right\rfloor<100.$ (26)
    According to this inequality, we obtain $t_{1}^{p_{a}}-t_{0}^{p_{a}}<900$. In order for a block $\mathcal{B}_{1}^{p_{a}}$ to be valid, we derive $t_{1}^{p_{a}}-t_{0}^{p_{a}}>0$. Since $t_{1}^{p_{a}}-t_{0}^{p_{a}}\in N$, we have:
    $t_{1}^{p_{a}}-t_{0}^{p_{a}}\in(1,900).$ (27)
    According to equations (25) and (27), Theorem 6 (1) is proved. Below we prove Theorem 6 (2).
  2. (2)
    Similarly, for $i=2$, we need to ensure that the following condition holds:
    $\mathcal{D}_{i}^{p_{a}}>\mathcal{D}_{i}^{p_{h}}\Rightarrow\mathcal{D}_{2}^{p_{a}}>\mathcal{D}_{2}^{p_{h}}.$ (28)
    According to the difficulty calculation formula, we have:
    $\mathcal{D}_{2}^{p_{h}}=\mathcal{D}_{1}^{p_{h}}+max\left\{1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor,-99\right\}\cdot\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor$ (29)
    and
    $\mathcal{D}_{2}^{p_{a}}=\mathcal{D}_{1}^{p_{a}}+max\left\{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99\right\}\cdot\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor.$ (30)
    Therefore, we can derive:
    $\mathcal{D}_{2}^{p_{a}}-\mathcal{D}_{2}^{p_{h}}=\overset{>0}{\overbrace{\left({\mathcal{D}_{1}^{p_{a}}-\mathcal{D}_{1}^{p_{h}}}\right)}}$ (31)
    $+ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99}\right\}}$
    $-max\left\{{1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor,-99}\right\}\geq 0.$
    We take the example of a minimum-risk attack (where the difficulty of an adversarial block at the same height in the SUUM model is one greater than that of an honest block).
    (2.1) For $\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor=\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor$, it suffices to ensure that the following condition holds:
    $ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99}\right\}}$ (32)
    $-max\left\{{1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor,-99}\right\}\geq 0.$
    Convert the above inequality into the following system of inequalities:
    $\left\{\begin{matrix}{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor\geq 1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor}&(1)\\ {1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor\geq-99}&(2)\end{matrix}\right.$. (33)
    To solve inequality (1), we have:
    $\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}-\left(t_{2}^{p_{a}}-t_{1}^{p_{a}}\right)}{9}\right\rfloor\geq 0.$ (34)
    To solve inequality (2), we have:
    $\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor\leq 100.$ (35)
    Therefore, the solution to the above system of inequalities is $\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}-\left(t_{2}^{p_{a}}-t_{1}^{p_{a}}\right)}{9}\right\rfloor\geq 0$ and $\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor\leq 100$.
    To ensure the legitimacy of block $\mathcal{B}_{2}^{p_{a}}$, the following conditions must be satisfied: $t_{2}^{p_{a}}-t_{1}^{p_{a}}>0$. Therefore, if $i=2$ and $\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor=\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor$, the SUUM attack is successful if and only if the following conditions hold: $\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}-\left(t_{2}^{p_{a}}-t_{1}^{p_{a}}\right)}{9}\right\rfloor\geq 0$ and $t_{2}^{p_{a}}-t_{1}^{p_{a}}\in[1,900)$.
    (2.2) For $\frac{\mathcal{D}_{1}^{p_{h}}+1}{2048}=\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor$, we have $\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor=\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor+1$. At this point, it is only necessary to ensure that the following conditions hold:
    $\begin{matrix}&ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor,-99}\right\}}\cdot\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor\\ &-ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99}\right\}}\cdot\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor\geq 0\\ &\Rightarrow ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}}{9}\right\rfloor,-99}\right\}}\cdot\left({\left\lfloor\frac{\mathcal{D}_{1}^{p_{s}}}{2048}\right\rfloor-1}\right)\\ &-ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99}\right\}}\cdot\left\lfloor\frac{\mathcal{D}_{1}^{p_{a}}}{2048}\right\rfloor\geq 0.\end{matrix}$
    Since $ma{x\left\{{1-\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor,-99}\right\}}<0$, it follows that:
[…]
    To solve inequality (1), we have:
    $\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}-\left({t_{2}^{p_{h}}-t_{1}^{p_{h}}}\right)}{9}\right\rfloor\geq 0.$ (37)
    To slove inequality (2), we have:
    $\left\lfloor\frac{t_{2}^{p_{a}}-t_{1}^{p_{a}}}{9}\right\rfloor\leq 100.$ (38)
    In order for a block $\mathcal{B}_{2}^{p_{a}}$ to be valid, it needs to be satisfied: $t_{2}^{p_{a}}-t_{1}^{p_{a}}\geq 1$. Thus we have:
    $t_{2}^{p_{a}}-t_{1}^{p_{a}}\in[1,900).$ (39)
    Combining inequalities (37) and (39), if $i=2$ and $\frac{\mathcal{D}_{1}^{p_{h}}+1}{2048}=\left\lfloor\frac{\mathcal{D}_{1}^{p_{h}}}{2048}\right\rfloor$, SUUM attack succeeds if and only if the following conditions hold: $\left\lfloor\frac{t_{2}^{p_{h}}-t_{1}^{p_{h}}-\left(t_{2}^{p_{a}}-t_{1}^{p_{a}}\right)}{9}\right\rfloor\geq 0$ and $t_{2}^{p_{a}}-t_{1}^{p_{a}}\in[1,900)$.
  3. (3)
    For $i\geq 3$, the proof process is analogous to the case where $i=2$.
Appendix H Proof of Theorem 7
To minimize the risk posed by the SUUM attack, which corresponds to minimizing the block difficulty growth rate, according to Theorem 6, we proceed as follows:
  1. (1)
    For $i=1$, we take the minimum value within the condition $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{a}}}{9}\right\rfloor\in[1,+\infty)$, i.e., $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{a}}}{9}\right\rfloor=1$.
  2. (2)
    For $i\geq 2$, we take the minimum value within the condition $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}-\left(t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\right)}{9}\right\rfloor\in[1,+\infty)$, i.e., $\left\lfloor\frac{t_{i}^{p_{h}}-t_{i-1}^{p_{h}}-\left(t_{i}^{p_{a}}-t_{i-1}^{p_{a}}\right)}{9}\right\rfloor=1$.
By integrating these two cases, we can prove the theorem.
Appendix I Proof of Theorem 8
Based on the analysis of state transitions in the Deployment State, Downgrade State, and Attack State of SUUM presented in Section V-B, we calculate the absolute and relative shares of coin-base rewards for both the SUUM adversary and honest participants.
The absolute share of coin-base rewards for the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ is:
$\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{SUUM}$ $=\mathbb{P}\left({Deploy}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$ (40)
$+\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$+\mathbb{P}\left({Attack~~i,i\geq 1}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\begin{pmatrix}{P\left({Deploy}\right)+P\left({Downgrade}\right)}\\ {+P\left({Attack~~i,i\geq 1}\right)}\end{pmatrix}\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathbb{P}_{\mathcal{A}}\left({any~t}\right).$
Therefore, the absolute share of coin-base rewards for the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ remains unchanged.
The absolute share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ is:
$\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{SUUM}$ $=\mathbb{P}\left({Deploy}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({t<9}\right)$ (41)
$+\mathbb{P}\left({Deploy}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)$
$+\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t<9}\right)$
$+\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{H}}\left({t\geq 9}\right)$
$+\mathbb{P}\left({Attack~~i+1,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)$
$-\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack~~i,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)$
$=\begin{pmatrix}{P\left({Deploy}\right)+P\left({Downgrade}\right)}\\ {+P\left({Attack~~i,i\geq 1}\right)}\end{pmatrix}\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack~~1}\right)\cdot\mathcal{P}_{\mathcal{H}}\left({any~t}\right)$
$-\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack~~i,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)$
$=\mathcal{P}_{\mathcal{H}}\left({any~t}\right)-\mathbb{P}\left({Attack~~1}\right)$
$-\mathbb{P}\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)$
$-\mathbb{P}\left({Attack~~i+1,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)$
$<\mathcal{P}_{\mathcal{H}}\left({any~t}\right).$
Therefore, the absolute share of coin-base rewards for honest participants decreases.
The relative share of coin-base rewards for the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ is:
$\mathbb{E}\left[\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{SUUM}\right]=\frac{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{SUUM}}{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{SUUM}+\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{SUUM}}$ (42)
$=\frac{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}{\begin{pmatrix}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)-P\left({Attack~~1}\right)}\\ {-P\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}\\ {+P\left({Attack~~i+1,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)}\end{pmatrix}}$
$>\frac{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}$
$=\mathbb{P}_{\mathcal{A}}\left({any~t}\right).$
Therefore, the expected relative share of coin-base rewards for the SUUM adversary $\mathcal{P}_{\mathcal{A}}$ increases.
The expected relative share of coin-base rewards for honest participant $\mathcal{P}_{\mathcal{H}}$ is:
$\mathbb{E}\left[\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{SUUM}\right]=\frac{\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{SUUM}}{\mathbb{R}_{\mathcal{P}_{\mathcal{A}}}^{SUUM}+\mathbb{R}_{\mathcal{P}_{\mathcal{H}}}^{SUUM}}$ (43)
$=\frac{\begin{pmatrix}{\mathbb{P}_{\mathcal{H}}\left({any~t}\right)-P\left({Attack~~1}\right)}\\ {-P\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}\\ {+P\left({Attack~~i+1,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)}\end{pmatrix}}{\begin{pmatrix}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)-P\left({Attack~~1}\right)}\\ {-P\left({Downgrade}\right)\cdot\mathbb{P}_{\mathcal{A}}\left({any~t}\right)}\\ {+P\left({Attack~~i+1,i\geq 1}\right)\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)}\end{pmatrix}}$
$<\frac{\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}{\mathbb{P}_{\mathcal{A}}\left({any~t}\right)+\mathcal{P}_{\mathcal{H}}\left({any~t}\right)}$
$=\mathcal{P}_{\mathcal{H}}\left({any~t}\right).$
Therefore, the expected relative share of coin-base rewards for honest participants $\mathcal{P}_{\mathcal{H}}$ decreases.
Appendix J Proof of Theorem 9
Firstly, it has been proved in the literature [26] that RUM outperforms honest mining. This provides us with a starting point for our subsequent comparisons.
Next, let’s focus on the comparison between UUM and RUM. When we look at their state space and transition processes, we can find that there is a key difference in the conditions that trigger the transition from the deployment state to the attack state.
For RUM, it transitions from the deployment state to the attack state when an honest participant finds a new block whose timestamp differs from the timestamp of the previous block in the main chain by a certain value. Specifically, this value needs to be greater than or equal to $9$ and less than $18$, which we can denote as $\mathcal{P}_{\mathcal{H}}\left({9\leq t<18}\right)$. In a practical blockchain scenario, this means that only when the time interval between the generation of consecutive blocks by honest participants falls within this specific range will RUM enter the attack state.
In contrast, for UUM, it transitions from the deployment state to the attack state when an honest participant finds a new block whose timestamp differs from the previous block’s timestamp in the main chain by a value greater than or equal to $9$, denoted as $\mathcal{P}_{\mathcal{H}}\left({t\geq 9}\right)$. For example, if we imagine the blockchain as a sequence of events over time, UUM is more likely to start an attack as it has a broader range of time differences that can trigger the attack state compared to RUM.
This difference in the transition conditions leads to a higher steady-state probability of being in the attack state for UUM compared to RUM. And as a result, UUM adversaries can obtain higher rewards than RUM adversaries.
Similarly, when we consider the comparison between SUUM and UUM, the approach to proving that SUUM outperforms UUM follows a similar logic. SUUM expands the adversary’s state space in a significant way. In the attacking state, adversaries in SUUM can release held-back blocks with carefully selected timestamps. Let’s say, they can choose the most opportune moments to release these blocks to maximize their impact on the blockchain’s operation and ultimately obtain higher rewards than UUM adversaries. This is somewhat like having more strategic options in a game, which gives SUUM an advantage over UUM.
Appendix K Proof of Theorem 10
We discuss the difficulty risks posed by each of the three types of attacks as follows:
  1. (1)
    First, we analyze the difficulty risk posed by the RUM attack. Recalling Theorems 2 and 3 from Section III, we understand that each successful execution of a RUM attack increases the difficulty risk of the blockchain by $1$. We use the notation $Attack\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy$ to represent this situation. That is, every time a RUM adversary successfully executes an attack and triggers the state change from the attack state to the deployment state, it leads to an increment in the blockchain’s difficulty risk. Consequently, the increased risk associated with the RUM attack can be quantified as the number of successful attacks by a RUM adversary, denoted by $\sum\left({Attack\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy}\right)$.
  2. (2)
    The proof of UUM follows a methodology analogous to that of RUM.
  3. (3)
    Finally, we analyze the difficulty risk posed by SUUM attack. Recalling Theorems 6 and 7 from Section V, we understand that each successful execution of a SUUM attack also increases the difficulty risk of the blockchain by $1$. This increase in risk is associated with two specific types of state transitions, represented by $Downgrade\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy$ and $Attack~1\overset{\mathcal{P}_{\mathcal{H}}~~any~t}{\Rightarrow}Deploy$. In other words, when an SUUM adversary successfully conducts an attack and makes the state change happen in either of these two ways, it adds to the overall risk related to the blockchain’s difficulty. Therefore, the increased risk associated with SUUM attack can be quantified as the number of successful attacks by a SUUM adversary, denoted by $\sum\begin{pmatrix}{Downgrade\overset{\mathcal{P}_{\mathcal{A}}~~any~t}{\Rightarrow}Deploy}\\ {+Attack~1\overset{\mathcal{P}_{\mathcal{H}}~~any~t}{\Rightarrow}Deploy}\end{pmatrix}$.
Appendix L Proof of Theorem 11
We discuss the forking rate posed by each of the three types of attacks and honest mining as follows:
  1. (1)
    For honest mining, all participants adhere to the protocol and engage in honest mining practices. Under such circumstances, no participant undertakes malicious actions deliberately intended to cause forks. Consequently, the deliberate forking rate induced by honest mining, denoted as ${\mathbb{F}\mathbb{R}}^{HM}=0$.
  2. (2)
    For the RUM attack, the adversary executes the RUM attack while other honest participants continue to mine honestly. Under such circumstances, revisiting the RUM attack methodology outlined in Section III, we understand that the deliberate forking of the blockchain occurs as a result of a successful RUM attack by the adversary. A successful RUM attack is contingent upon the RUM adversary, who is in the attack state, successfully generating the next block, with the timestamp difference between it and its parent block being less than $9$. Therefore, the deliberate forking rate induced by the RUM attack, denoted as ${\mathbb{F}\mathbb{R}}^{RUM}=\mathbb{P}_{Attack}^{RUM}\cdot\mathcal{P}_{\mathcal{A}}(t<9)$.
  3. (3)
    For the UUM attack, the adversary executes the UUM attack while other honest participants continue to mine honestly. Under such circumstances, revisiting the UUM attack methodology outlined in Section IV, we understand that the deliberate forking of the blockchain arises due to a successful UUM attack by the adversary. A successful UUM attack occurs if and only if the UUM adversary, who is in the attack state, successfully generates the next block and strategically manipulates its timestamp such that the adversary’s block in the fork competition is preferentially selected by other honest participants over the honest chain. Therefore, the deliberate forking rate induced by the UUM attack, denoted as ${\mathbb{F}\mathbb{R}}^{UUM}=\mathbb{P}_{Attack}^{UUM}\cdot\mathcal{P}_{\mathcal{A}}(any~t)$.
  4. (4)
    For SUUM attack, the adversary executes the SUUM attack while other honest participants continue to mine honestly. In such a scenario, revisiting the SUUM attack methodology outlined in Section V, we can identify two scenarios that lead to deliberate forking of the blockchain. The first scenario occurs when the blockchain topology is in the attack state, where the withholding of blocks by the SUUM adversary will inevitably cause a blockchain fork. The corresponding probability for this scenario is $\sum_{i=0}^{n}\mathbb{P}_{Attack~i}^{SUUM}$. The second scenario is when the SUUM adversary, who is in the downgrade state, successfully generates the next block and strategically manipulates its timestamp such that this block is preferentially selected, leading to deliberate forking of the blockchain. The corresponding probability for this scenario is $\mathbb{P}_{Downgrade}^{SUUM}\cdot\mathcal{P}_{\mathcal{A}}(any~t)$. Therefore, the deliberate forking rate induced by the SUUM attack, denoted as ${\mathbb{F}\mathbb{R}}^{SUUM}$, is given by the sum of the probabilities of these two scenarios:
    ${\mathbb{F}\mathbb{R}}^{SUUM}$ $=\sum_{i=0}^{n}\mathbb{P}_{Attack~i}^{SUUM}+\mathbb{P}_{Downgrade}^{SUUM}\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)$ (44)
    $=\mathbb{P}_{Attack}^{SUUM}+\mathbb{P}_{Downgrade}^{SUUM}\cdot\mathcal{P}_{\mathcal{A}}(any~t).$
Next, we proceed to compare the magnitudes of the forking rates arising from these four scenarios. On the one hand, according to Theorem 9, for adversaries possessing equivalent computational power, the steady-state probabilities of being in the attack state for SUUM, UUM, and RUM decrease sequentially, i.e.,
$\mathbb{P}_{Attack}^{SUUM}>\mathbb{P}_{Attack}^{UUM}>\mathbb{P}_{Attack}^{RUM}.$ (45)
On the other hand, it is evident that
$1>\mathcal{P}_{\mathcal{A}}\left({any~t}\right)>\mathcal{P}_{\mathcal{A}}(t<9).$ (46)
Therefore, we conclude that
$\mathbb{P}_{Attack}^{SUUM}>\mathbb{P}_{Attack}^{UUM}\cdot\mathcal{P}_{\mathcal{A}}\left({any~t}\right)>\mathbb{P}_{Attack}^{RUM}\cdot\mathcal{P}_{\mathcal{A}}\left({t<9}\right)$ (47)
$\Rightarrow$ ${\mathbb{F}\mathbb{R}}^{SUUM}>{\mathbb{F}\mathbb{R}}^{UUM}>{\mathbb{F}\mathbb{R}}^{RUM}>{\mathbb{F}\mathbb{R}}^{HM}.$
Appendix M Proof of Theorem 12
We know that the relationship between targettarget and the difficulty $\mathcal{D}$ can be expressed as:
$target=\frac{max\_target}{\mathcal{D}}.$ (48)
Where, max_targetmax\_target is a constant, and for Ethereum 1.x, its value is $2^{256}$.
The probability $P_{\mathcal{A}}$ for adversary $\mathcal{A}$ to find the next block can be calculated as follows:
$\mathcal{P}_{\mathcal{A}}=\mu_{\mathcal{A}}\cdot\frac{1}{target}.$ (49)
Since targettarget is inversely proportional to the difficulty $\mathcal{D}$, we can replace target with an expression related to $\mathcal{D}$:
$\mathcal{P}_{\mathcal{A}}=\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}}{max\_target}.$ (50)
  1. 1.
    For RUM attack, its cost can be expressed as:
    ${\mathbb{A}\mathbb{C}}^{RUM}$ $=\mathcal{P}_{\mathcal{A}}^{\mathcal{B}_{0}^{p_{h}}}-\mathcal{P}_{\mathcal{A}}^{\mathcal{B}_{1}^{p_{h}}}$ (51)
    $=\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}_{0}^{p_{h}}}{max\_target}-\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}_{1}^{p_{h}}}{max\_target}$
    $=\mu_{\mathcal{A}}\cdot\frac{\left({\mathcal{D}_{0}^{p_{h}}-\mathcal{D}_{1}^{p_{h}}}\right)}{max\_target}.$
    According to [26], the initiation condition for the RUM attack is $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor\in[1,9)$, i.e., $\mathcal{D}_{0}^{p_{h}}=\mathcal{D}_{1}^{p_{h}}$. Therefore, we have:
    ${\mathbb{A}\mathbb{C}}^{RUM}=0.$ (52)
  2. 2.
    For UUM and/or SUUM attacks, their cost can be expressed as:
    ${\mathbb{A}\mathbb{C}}^{UUM}$ $={\mathbb{A}\mathbb{C}}^{SUUM}=\mathcal{P}_{\mathcal{A}}^{\mathcal{B}_{0}^{p_{h}}}-\mathcal{P}_{\mathcal{A}}^{\mathcal{B}_{1}^{p_{h}}}$ (53)
    $=\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}_{0}^{p_{h}}}{max\_target}-\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}_{1}^{p_{h}}}{max\_target}$
    $=\mu_{\mathcal{A}}\cdot\frac{\left({\mathcal{D}_{0}^{p_{h}}-\mathcal{D}_{1}^{p_{h}}}\right)}{max\_target}.$
    According to Theorem 1 and Theorem 6, the initiation condition for UUM and SUUM attacks is $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor\in[9,+\infty)$, i.e., $\mathcal{D}_{0}^{p_{h}}-\mathcal{D}_{1}^{p_{h}}\geq 1$. Therefore, we have:
    ${\mathbb{A}\mathbb{C}}^{UUM}={\mathbb{A}\mathbb{C}}^{SUUM}$ $=\mu_{\mathcal{A}}\cdot\frac{\left({\mathcal{D}_{0}^{p_{h}}-\mathcal{D}_{1}^{p_{h}}}\right)}{max\_target}$ (54)
    $>\mu_{\mathcal{A}}\cdot\frac{1}{max\_target}$
    $>0.$
    Taking Ethereum 1.x as an example, by substituting max_targetmax\_target with $2^{256}$, we obtain:
    $\mathcal{P}_{\mathcal{A}}=\mu_{\mathcal{A}}\cdot\frac{\mathcal{D}}{2^{256}}.$ (55)
    Thus, the probability $\mathcal{P}_{\mathcal{A}}$ for an adversary $\mathcal{A}$, possessing a power ratio of $\mu_{\mathcal{A}}$, to find the next block is the proportion of its power to the total power, multiplied by the ratio of the block difficulty $\mathcal{D}$ to the maximum target value.
    It is noteworthy that, upon reviewing historical data from Ethereum 1.x, we observe that the maximum difference in difficulty between two consecutive blocks is approximately $8.86441324\times 10^{12}\approx 2^{39.65}$. Therefore, we can derive:
    ${\mathbb{A}\mathbb{C}}^{UUM}={\mathbb{A}\mathbb{C}}^{SUUM}$ $\leq\mu_{\mathcal{A}}\cdot\frac{max\left\{{\mathcal{D}_{0}^{p_{h}}-\mathcal{D}_{1}^{p_{h}}}\right\}}{2^{256}}$ (56)
    $\leq\mu_{\mathcal{A}}\cdot\frac{2^{39.65}}{2^{256}}\approx\mu_{\mathcal{A}}\cdot\frac{1}{2^{216.35}}$
    $\approx 0.$
Appendix N Mitigations
N-A Adjust the Difficulty Adjustment Algorithm
The difficulty adjustment algorithm of Ethereum 1.x is at risk of being exploited in the face of Uncle Maker attacks. Adversaries can manipulate the timestamp to influence the block difficulty and thus gain an unfair advantage. Therefore, it is necessary to design a more robust and manipulation-resistant difficulty adjustment algorithm. For example, more factors can be considered to determine the block difficulty. Instead of relying solely on the timestamp and the difficulty of the parent block, dynamic factors such as the actual computational power distribution of the network and the number of transactions can also be incorporated. In this way, it becomes difficult for adversaries to simply manipulate the timestamp to reduce their attack difficulty, thereby increasing the cost and difficulty of the attack and reducing the probability of the attack occurring.
One possible improvement direction is to adopt a moving average-based difficulty adjustment method, which takes into account the generation time and difficulty of multiple past blocks instead of just the current block and its parent block. This can smooth the difficulty adjustment process and reduce the exploitability of short-term fluctuations by adversaries. Additionally, upper and lower limits for difficulty adjustment can be set to prevent adversaries from causing significant drops or rises in difficulty through extreme timestamp manipulations, thereby maintaining the stability and predictability of the blockchain difficulty.
N-B Strengthen Timestamp Verification
Given that the Uncle Maker attack is highly reliant on the manipulation of timestamps, strengthening the timestamp verification mechanism is one of the key measures to mitigate such attacks. A more rigorous and reliable source of timestamps can be established. For instance, a network of distributed timestamp servers can be employed to ensure that all nodes can obtain consistent and accurate time information. Simultaneously, stricter limitations on the range of block timestamps should be imposed to prevent adversaries from arbitrarily setting unreasonable timestamps. For example, it can be stipulated that the block timestamp must fall within a reasonable interval, be in line with the average block generation time of the network, and possess reasonable continuity with the timestamps of preceding blocks.
A timestamp consensus algorithm can be introduced, mandating that multiple nodes perform consensus verification on the block timestamp. Only when a certain proportion of mining power approve can the timestamp be regarded as valid. Additionally, for blocks with abnormal timestamps, more in-depth examination and verification can be carried out. Even the block can be temporarily rejected until the legality of its timestamp is ascertained. This can effectively prevent adversaries from launching Uncle Maker attacks by manipulating timestamps and enhance the security of the blockchain system.
N-C Introduce an Economic Penalty Mechanism
To further deter adversaries, an economic penalty mechanism can be introduced. For nodes or participants detected to be involved in the Uncle Maker attack, in addition to confiscating the improper gains obtained through the attack, additional economic penalties can be imposed on them, such as deducting a certain proportion of collateral assets (if applicable) or future mining rewards. This can increase the cost of the attack and make adversaries more cautious when considering launching an attack.
A reporting mechanism can be established to encourage other nodes to report suspicious attack behaviors. For nodes that report successfully, a certain reward can be given, and the source of the reward can be the fines imposed on the adversaries. In this way, a community supervision mechanism can be formed to jointly maintain the security and fairness of the blockchain. At the same time, the attack behaviors and penalty situations should be regularly publicized to serve as a warning and reduce the number of potential adversaries.
N-D Improve Network Monitoring and Early Warning Capabilities
Real-time monitoring of the operating status of the blockchain network and timely detection of abnormal block generation patterns and timestamp behaviors are crucial for quickly responding to and mitigating Uncle Maker attacks. Specialized monitoring tools and algorithms can be deployed to perform real-time monitoring and analysis of key indicators such as block timestamps, difficulty changes, and uncle block generation in the network. Once an abnormal situation is detected, an early warning signal should be issued in a timely manner to notify network participants to take corresponding measures, such as suspending the acceptance of suspicious blocks and activating the emergency response mechanism.
Utilize machine learning and artificial intelligence technologies to conduct in-depth analysis of network data and establish prediction models for attack behaviors. Through learning from historical data, patterns and trends that may indicate the imminent occurrence of Uncle Maker attacks can be identified, and preventive measures can be taken in advance. For example, if it is found that the block timestamp setting pattern of a certain node or participant is significantly different from normal behavior and is accompanied by an abnormal increase in the frequency of uncle block generation, the system can automatically mark the node as a suspicious object and strengthen the monitoring of its subsequent behaviors. At the same time, the detected abnormal situations should be shared with the entire blockchain community to promote the community to jointly respond to attack threats and improve the security and stability of the entire network.